### v4

# Scale Reliant Inference

Abstract
Scientific fields such as genomics, ecology, and political science often collect multivariate count data. In these fields, the data are often sufficiently noisy such that inferences regarding the total size of the measured systems have substantial uncertainty. This uncertainty can hinder downstream analyses, such as differential analysis in case-control studies. There have historically been two approaches to this problem: one considers the data as compositional and the other as counts that can be normalized. In this article, we use the framework of partially identified models to rigorously study the types of scientific questions (estimands) that can be answered (estimated) using these data. We prove that satisfying Frequentist inferential criteria is impossible for many estimation problems. In contrast, we find that the criteria for Bayesian inference can be satisfied, yet it requires a particular type of model called a Bayesian partially identified model. We introduce Scale Simulation Random Variables as a flexible and computationally efficient form of Bayesian partially identified models for analyzing these data. We use simulations and data analysis to validate our theory.
Keywords: Multivariate Analysis, Identifiability, Uncertainty Quantification, Sequence Count Data
1 Introduction
Many scientific fields collect noisy multivariate measurements where the scale of the measurements (i.e., total measured amount) is arbitrary and does not represent the scale of the underlying system (i.e., true total amount). These data occur frequently in genomics (Lovell et al.,, 2011; Gloor et al.,, 2017), macro-ecology (Murphy et al.,, 2016; Jackson,, 1997), epidemiology (Silverman et al.,, 2020), demography (Rossi et al.,, 1987), and political polling (Topîrceanu and Precup,, 2020). Consider forecasting a national election based on a collection of polls conducted in distinct sub-national regions. The scale of the polls may be incomparable such that a single vote in one poll represents a different number of total votes than a single vote in a different poll. The observed counts could be normalized for downstream analysis using survey weighting if accurate information on the size of the voter roll is available for each region. However, the problem is substantially more complicated if such data is lacking or subject to errors. Similar challenges are faced when studying gut microbiota, where fecal samples are sequenced to quantify the abundance of different microbes in each subject’s gut. Prior work has established that the measurement scale (e.g., the total number of sequenced DNA molecules) is driven by technical factors: it is unrepresentative of meaningful biological variation in total microbial load (Lovell et al.,, 2011; Props et al.,, 2017; Vandeputte et al.,, 2017; Paulson et al.,, 2013). Statisticians often agree that these issues of arbitrary data scale can confound analyses (Topîrceanu and Precup,, 2020). For example, in studying microbiota, it is well known that methods used to account for this arbitrary scale can dominate study conclusions (Roche and Mukherjee,, 2022; Weiss et al.,, 2017). However, there are fundamental disagreements regarding how to address these challenges and what types of analyses can be rigorously performed (Gloor et al.,, 2017; Lovell et al.,, 2015, 2011; Gloor et al.,, 2017). This paper addresses the question: when does the arbitrary scale of the data limit the types of analyses that can be accurately carried out and which methods are best suited for mitigating undesirable inference properties such as bias and Type-I error?
Some authors call these data compositional (Gloor et al.,, 2017). This term reflects the idea that if the data scale is arbitrary, then the data only contains information about the system’s composition (i.e., proportions). According to the Scale Invariance Principle of Compositional Data Analysis (CoDA), if the observed data are compositional, then analyses must be scale invariant: they must focus solely on the relative frequency of the counts (Greenacre et al.,, 2023; Pawlowsky-Glahn et al.,, 2015). For example, within the study of human microbiota, proponents of the CoDA perspective have argued that inter-taxa correlations cannot be estimated rigorously as they are a function of microbial load and cannot be expressed in terms of microbial compositions alone (Lovell et al.,, 2015). Instead, CoDA proponents have suggested that authors focus their attention on scale invariant statistics such as proportionality (Lovell et al.,, 2015). In short, the field of CoDA develops tools to answer scale invariant research questions and often dismisses scale reliant questions as unanswerable (Aitchison,, 1982; Lovell et al.,, 2015, 2011).
CoDA’s scale invariant perspective is rigorous, but the rigor comes with constraints on the scientific question. While constraining ourselves to scale invariant questions seems logical when we cannot estimate the system scale accurately, it is sub-optimal in less extreme settings. In the setting of polling, additional measurements of scale (e.g., the size of the voter roll) could be used to overcome these limits and rigorously answer scale reliant questions. Alternatively, even if we lack scale measurements, we may have prior knowledge that allows us to quantify our uncertainty about the system scale. This paper formally defines the conditions under which we can move beyond CoDA’s insistence on scale invariance and answer scale reliant questions using only incomplete information about the system scale.
While this paper is the first to formalize the conditions under which scale reliant questions can be answered using data where the scale/total counts are noisy or missing, many models, algorithms, and methods have been proposed to address common scale reliant inferential tasks such as correlation analysis or differential analysis (Lovell et al.,, 2015; Fernandes et al.,, 2014; Love et al.,, 2014). These include data normalizations (Weiss et al.,, 2017), categorical data models with identification restrictions (Rabe-Hesketh and Skrondal,, 2001), or external measurements of scale (Gibson and Gerber,, 2018; Williamson et al.,, 2019), as well as a wide variety of methods that assume sparsity in the underlying system (Grantham et al.,, 2020). There are also many other approaches, and we cannot review them all. Instead, motivated by recent debates within the genomics literature (e.g.,  Gloor et al., (2017)), we bias our presentation towards examples from that field to make the theory in this paper concrete.
The main contributions of this paper include the following. Using the concept of partial identifiability (Gustafson,, 2015), we state the conditions under which the estimation of estimands of scientific interest, such as correlations between variables or differential analysis between case versus control, are possible and reliable. We show that Frequentist procedures do not provide reliable estimation of these estimands. Then, we demonstrate that a particular class of Bayesian models called Bayesian Partially Identified Models (Bayesian PIMs) can result in rigorous inference. We also provide a simulation framework called Scale Simulation Random Variables (SSRVs), which can simplify the specification and inference of Bayesian PIMs. We show that SSRVs allow researchers to represent more realistic beliefs about scale than is possible with current methods.
2 Motivating Example
The analysis of gut microbiota data often centers around a $D\times N$ sequence count table $\mathbf{Y}$ with elements $\mathbf{Y}_{dn}$ denoting the number of sequenced (e.g., measured) DNA molecules that map to a particular microbial taxon $d$ in fecal sample $n$. The measurement process is complex and requires many experimental and computational steps to transform fecal samples into a count table $\mathbf{Y}$ (Caporaso et al.,, 2011). Researchers frequently analyze these data in the hopes of identifying which, if any, taxa change in abundance between cases and controls. This is referred to as differential abundance analysis. As a motivating example, we will discuss the complexity of this problem and later show that standard methods for performing differential abundance analysis can be very inaccurate.
One can consider the measurement process producing the sequence count table $\mathbf{Y}$ as a type of poll: there is a pool of microbes that could be measured, but only an arbitrary sized subset of those are actually collected (Fernandes et al.,, 2014; Lovell et al.,, 2011). Let $\mathbf{W}$ denote a $D\times N$ matrix with elements $\mathbf{W}_{dn}$ representing the true amount of taxon $d$ in the gut from which fecal sample $n$ was obtained. Note that this matrix can be uniquely described in terms of its composition $\mathbf{W}^{{\tiny\mbox{comp}}}$ and scale $W^{{\tiny\mbox{tot}}}$ via:
$\mathbf{W}_{dn}$ $=\mathbf{W}^{{\tiny\mbox{comp}}}_{dn}\,\,W^{{\tiny\mbox{tot}}}_{n}$
$W^{{\tiny\mbox{tot}}}_{n}$ $=\sum_{d=1}^{D}\mathbf{W}_{dn}.$
This implies that $\mathbf{W}^{{\tiny\mbox{comp}}}$ is a $D\times N$ matrix with columns that sum to $1$, and $W^{{\tiny\mbox{tot}}}$ is a positive-valued $N$-vector. When we say that the scale of the data is arbitrary, we mean that, even asymptotically, $\mathbf{Y}$ does not contain sufficient information to uniquely identify the system scale $W^{{\tiny\mbox{tot}}}$. This statement is formally defined in Section 4.
One of the most common analyses of microbiome data is called Differential Abundance Analysis (DA). In DA, there is a binary covariate ($x_{n}\in\{0,1\}$) associated with each measurement which denotes one of two mutually exclusive conditions (e.g., case versus control). The simplest estimand for DA is the log-fold-change (LFC). The LFC of taxon $d$, $\theta_{d}$, is often defined as the difference in means of log-transformed abundances (Love et al.,, 2014):
$\theta_{d}=\operatorname*{mean}_{n:x_{n}=1}\log\mathbf{W}_{dn}-\operatorname*{mean}_{n:x_{n}=0}\log\mathbf{W}_{dn}.$ (1)
Estimating LFCs requires knowledge of the system scale $W^{{\tiny\mbox{tot}}}$. Based on the decomposition of $\mathbf{W}_{dn}$, we can write $\log\mathbf{W}_{dn}=\log\mathbf{W}^{{\tiny\mbox{comp}}}_{dn}+\log W^{{\tiny\mbox{tot}}}_{n}$, thus
$\theta_{d}$ $=\underbrace{\operatorname*{mean}_{n:x_{n}=1}(\log\mathbf{W}^{{\tiny\mbox{comp}}}_{dn})-\operatorname*{mean}_{n:x_{n}=0}(\log\mathbf{W}^{{\tiny\mbox{comp}}}_{dn})}_{\theta^{\tiny\mbox{comp}}_{d}}+\underbrace{\operatorname*{mean}_{n:x_{n}=1}(\log W^{{\tiny\mbox{tot}}}_{n})-\operatorname*{mean}_{n:x_{n}=0}(\log W^{{\tiny\mbox{tot}}}_{n})}_{\theta^{\tiny\mbox{tot}}}$
$=\theta^{\tiny\mbox{comp}}_{d}+\theta^{\tiny\mbox{tot}}.$ (2)
The main obstacle in obtaining accurate estimates of this estimand is that the scale of the data $\mathbf{Y}$ is typically arbitrary, and, as a result, we lack information about $W^{{\tiny\mbox{tot}}}$ and, therefore, $\theta^{\tiny\mbox{tot}}$.
To illustrate the problem, suppose that a researcher concludes that taxon $d$ is increased in cases as opposed to controls ($\theta_{d}>0$) based on the observation that the proportional abundance of taxon $d$ increases ($\theta^{\tiny\mbox{comp}}_{d}>0$). This conclusion is not supported by the observed data as knowledge of $\theta^{\tiny\mbox{comp}}_{d}$ is insufficient to determine $\theta_{d}$. Suppose the mean scale in the case condition is lower than the mean scale in the control condition ($\theta^{\tiny\mbox{tot}}<0$). Then, $\theta^{\tiny\mbox{comp}}_{d}$ will be spuriously large. In fact, if $|\theta^{\tiny\mbox{tot}}|>|\theta^{\tiny\mbox{comp}}_{d}|$, then $\theta^{\tiny\mbox{comp}}_{d}$ suggests that taxon $d$ is increasing when it is actually decreasing.
In Supplementary Section A, we provide a toy example with just $D=4$ taxa so that readers can walk through calculations themselves and see the spurious conclusions that can arise when analyzing data with arbitrary scale. Later, in Section 7, we provide a larger ($D=20$) simulation and an analysis of real data which demonstrates how two popular methods for DA analysis can demonstrate elevated Type I and/or Type-II error rates due to the problem of scale. For example, in Section 7.1, we demonstrate that two popular methods used for DA analysis have Type-I Error rates of over 70%. Contrary to standard statistical wisdom, we find these error rates actually increase with more data. Still, in those sections, we show that rigorous analyses are possible when statistical methods account for potential errors in modeling assumptions about scale.
3 Notation
We use $P_{a,b}$ to denote a probability model which is a function mapping from a parameter space $\mathcal{A}\times\mathcal{B}$ to a set of distributions $\mathcal{P}$. For a random variable $x$ which can be described by this model, we use the notation $P_{a,b}(x)$ to denote specific distribution of $x$ within the set $\mathcal{P}$ (i.e., a point within the image of the model $P_{a,b}$). The identification region ${\mbox{Id}}_{a,b}(P)$ is the pre-image of the model at a specific distribution $P\in\mathcal{P}$:
${\mbox{Id}}_{a,b}(P)=\{(a,b)\in\mathcal{A}\times\mathcal{B}:P_{a,b}=P\}.$
The model is identified at $P$ if ${\mbox{Id}}_{a,b}(P)$ is a singleton set. We define the marginal identification region of $a$ as
${\mbox{Id}}_{a}(P)=\{a\in\mathcal{A}:\exists\>b\in\mathcal{B},P_{a,b}=P\}.$
The marginal identification region of $b$ (${\mbox{Id}}_{b}(P)$) is defined analogously. Statements about the identifiability of a single parameter within a multi-parameter model refer to the marginal identification region: i.e., $a$ is identified at $P$ if $Id_{a}(P)$ is a singleton set. While identification can depend on the specific point $P\in\mathcal{P}$, for brevity, we will often omit this detail when the specific point can be inferred from context.
4 Formal Problem Statement
Here, we formally define what it means to ask a scale reliant question using data with arbitrary scale. As mentioned, we call this the problem of Scale Reliant Inference (SRI). First, we define the problem formally and then discuss how the problem statement relates to real-world problems such as DA analysis.
Consider the observed data $\mathbf{Y}$ as generated by a probability model $P_{\mathbf{W}^{{\tiny\mbox{comp}}},W^{{\tiny\mbox{tot}}}}$. We say that $\mathbf{Y}$ has arbitrary scale if the parameter $W^{{\tiny\mbox{tot}}}$ is unidentified at $P_{\mathbf{W}^{{\tiny\mbox{comp}}},W^{{\tiny\mbox{tot}}}}(\mathbf{Y})$. For simplicity, this article assumes that $\mathbf{W}^{{\tiny\mbox{comp}}}$ is identified. We make this assumption to focus our attention on the impact of scale and not other sources of measurement limitations associated with $\mathbf{W}^{{\tiny\mbox{comp}}}$ (e.g., PCR bias (Silverman et al.,, 2021)). We will denote data generation models where $\mathbf{W}^{{\tiny\mbox{comp}}}$ can be identified but $W^{{\tiny\mbox{tot}}}$ cannot as data with arbitrary scale or compositional surveys. We choose this name to highlight the connection to compositional data. However, the setting we consider is more general than the classic CoDA setting. We do not impose the simplex constraint of compositional data, and, therefore, we can define and model compositional surveys where each measurement, $\mathbf{Y}_{\cdot n}$, could be integer-valued and need not sum to a constant value.
The research questions we consider can be framed as learning the map $\mathbf{Y}\mapsto\theta$ where $\theta:=\theta\left(\mathbf{W}^{{\tiny\mbox{comp}}},W^{{\tiny\mbox{tot}}}\right)$. We call $\theta$ a target estimand because it is the quantity a researcher wants to estimate. We also introduce a nuisance function $\lambda=\lambda(\mathbf{W}^{{\tiny\mbox{comp}}},W^{{\tiny\mbox{tot}}})$ to ensure that the following concatenated function $f$ defined as
$(\theta,\lambda)=f(\mathbf{W}^{{\tiny\mbox{comp}}},W^{{\tiny\mbox{tot}}})=(\theta(\mathbf{W}^{{\tiny\mbox{comp}}},W^{{\tiny\mbox{tot}}}),\lambda(\mathbf{W}^{{\tiny\mbox{comp}}},W^{{\tiny\mbox{tot}}}))$
is invertible. With this condition, the model $P_{\mathbf{W}^{{\tiny\mbox{comp}}},W^{{\tiny\mbox{tot}}}}$ can be reparameterized in terms of $(\theta,\lambda)$ so $P_{\mathbf{W}^{{\tiny\mbox{comp}}},W^{{\tiny\mbox{tot}}}}=P_{f^{-1}(\theta,\lambda)}$. We define SRI in terms of the identifiability of $\theta$ within this reparameterized model.
Definition 1 (Scale Reliant and Scale Invariant Inference).
If $\mathbf{Y}$ is a compositional survey, the problem of estimating $\theta$ is Scale Reliant Inference (SRI) if $\theta$ is unidentified at $P_{\theta,\lambda}(\mathbf{Y})$. Otherwise (if $\theta$ is identified), the estimation problem is Scale Invariant Inference (SII).
To help motivate this definition, consider that, if $\mathbf{Y}$ is a compositional survey, the only way that $\theta$ can be unidentified is if it depends on $W^{{\tiny\mbox{tot}}}$ (since $\mathbf{Y}$ provides information about $\mathbf{W}^{{\tiny\mbox{comp}}}$ and $\theta$ is a known function). Therefore, our problem statement provides focus to our study: it restricts our attention to the setting where the lack of information about $W^{{\tiny\mbox{tot}}}$ contained in $\mathbf{Y}$ impedes our ability to uniquely identify $\theta$.
4.1 A Continuum between Scale Invariant and Scale Reliant Inference
Identifiability is not a dichotomous state but rather a continuum based on the identification region of a parameter. This implies that there is a spectrum between SRI and SII. Here, we define this progression before showing how these definitions apply to inferential problems described in existing literature.
In the extreme case where ${\mbox{Id}}_{\theta}(P)$ is a singleton set, $\theta$ is identified, and the estimation problem is SII. At the opposite end, where ${\mbox{Id}}_{\theta}(P)=\Theta$, $\theta$ is fully unidentified, and the estimation problem is Full SRI. Full SRI represents the extreme case where, due to the arbitrary scale of the data, all values of $\theta\in\Theta$ are consistent with the observed measurements. In between these two cases, where ${\mbox{Id}}_{\theta}(P)\subsetneq\Theta$, $\theta$ is partially identified, and the estimation problem is Partial SRI.
In Section 5, we show that the distinction between SII, Full SRI, and Partial SRI is directly related to the conditions under which certain types of inference (e.g., Frequentist versus Bayesian) are possible. Before providing those results, we first motivate that all three types of estimation problems are encountered within the literature. To do this, we describe the two relevant factors that determine whether a problem is SII, Full SRI, or Partial SRI: the degree to which the estimand depends on scale and the size of the identification region of $W^{{\tiny\mbox{tot}}}$ at $P_{\mathbf{W}^{{\tiny\mbox{comp}}},W^{{\tiny\mbox{tot}}}}(\mathbf{Y})$.
4.1.1 A Characterization of Target Estimands
Consider the following three target estimands:
$\eta_{n}$ $=\log\frac{\mathbf{W}_{d_{i}n}}{\mathbf{W}_{d_{j}n}},$ (3)
$\theta_{d}$ $=\operatorname*{mean}_{n:x_{n}=1}\log\mathbf{W}_{dn}-\operatorname*{mean}_{n:x_{n}=0}\log\mathbf{W}_{dn},$ (4)
$\Sigma_{d_{i},d_{j}}$ $=\text{cov}(\log\mathbf{W}_{d_{i}\cdot},\log\mathbf{W}_{d_{j},\cdot}).$ (5)
The first is the type of scale invariant quantity purported by the field of CoDA. To see why it does not depend on scale consider that
$\eta_{n}=\log\frac{\mathbf{W}_{d_{i}n}}{\mathbf{W}_{d_{j}n}}=\log\frac{\mathbf{W}^{{\tiny\mbox{comp}}}_{d_{i}n}}{\mathbf{W}^{{\tiny\mbox{comp}}}_{d_{j}n}}.$ (6)
The second is the LFC estimand introduced in Section 2 while the third is a quantity of central interest to micro- and macro-ecological modeling (e.g., Friedman and Alm, (2012)). While both have been criticized in the CoDA literature as not being scale invariant (Lovell et al.,, 2015), these estimands differ in the degree to which they rely on scale. Next we show that LFC estimand is fully scale reliant while the covariance estimand is partially scale reliant.
We define the scale invariance/reliance of an estimand as follows. Let $\mathbf{W}^{{\tiny\mbox{comp}}}\times\mathcal{W}^{\tiny\mbox{tot}}$ denote the set of all systems that have a particular composition $\mathbf{W}^{{\tiny\mbox{comp}}}$. Let $\theta(\mathbf{W}^{{\tiny\mbox{comp}}}\times\mathcal{W}^{\tiny\mbox{tot}}$) denote the image of this set with respect to the target estimand:
$\theta(\mathbf{W}^{{\tiny\mbox{comp}}}\times\mathcal{W}^{\tiny\mbox{tot}})=\{\theta\in\Theta:\exists\,\,W^{{\tiny\mbox{tot}}}\in\mathcal{W}^{\tiny\mbox{tot}},\theta(\mathbf{W}^{{\tiny\mbox{comp}}},W^{{\tiny\mbox{tot}}})=\theta\},$
where $\mathcal{W}^{\tiny\mbox{tot}}$ is the space of scale vectors (e.g., $\mathbb{Z}^{N+}$ if $W$ is integer valued or $\mathbb{R}^{N+}$ if continuous). We define scale invariance/reliance based on the size of this image. If $\theta(\mathbf{W}^{{\tiny\mbox{comp}}}\times\mathcal{W}^{\tiny\mbox{tot}})$ is a singleton set, we say that $\theta$ is scale invariant at $\mathbf{W}^{{\tiny\mbox{comp}}}$; otherwise, it is scale reliant at $\mathbf{W}^{{\tiny\mbox{comp}}}$. We sub-categorize scale reliant estimands as fully scale reliant if $\mbox{span}(\theta(\mathbf{W}^{{\tiny\mbox{comp}}}\times\mathcal{W}^{\tiny\mbox{tot}}))=\Theta$ or partially scale reliant if $\mbox{span}(\theta(\mathbf{W}^{{\tiny\mbox{comp}}}\times\mathcal{W}^{\tiny\mbox{tot}}))\subsetneq\Theta$. We will drop the phrase at $\mathbf{W}^{{\tiny\mbox{comp}}}$ if an estimand is scale invariant or reliant for all $\mathbf{W}^{{\tiny\mbox{comp}}}\in\mathcal{W}^{\tiny\mbox{comp}}$ where $\mathcal{W}^{\tiny\mbox{comp}}$ is the space of compositional matrices. To summarize, using only compositional information $\mathbf{W}^{{\tiny\mbox{comp}}}$, we can learn everything about scale invariant estimands, something about partially scale reliant estimands, and nothing about fully scale reliant estimands.
It follows from Equation 6 that the log-ratio estimand is scale invariant based on this definition. Moreover, from Equation 2, it follows that the LFC estimand is fully scale reliant. In Supplementary Section B.1, we prove that the covariance estimand is partially scale reliant so long as the covariance is finite. Elsewhere, we have shown how common differential set analyses (e.g., gene set enrichment analysis (Subramanian et al.,, 2005)) correspond to estimands that are scale reliant for some $\mathbf{W}^{{\tiny\mbox{comp}}}\in\mathcal{W}^{\tiny\mbox{comp}}$ and scale invariant for others (McGovern et al.,, 2023). In short, there is a tremendous diversity of estimands that goes beyond whether they are scale invariant or not.
4.1.2 Characterization of Compositional Surveys
Subtlety is required when considering compositional surveys. When authors equate data with arbitrary scale with compositional data (Gloor et al.,, 2017)), they are assuming that the observed data provides no information about the system scale. This corresponds to an extreme from of compositional survey where ${\mbox{Id}}_{W^{{\tiny\mbox{tot}}}}(P)=\mathcal{W}^{\tiny\mbox{tot}}$. We call these full compositional surveys. In some cases (e.g., human gut microbiome surveys), this may be a reasonable model. Yet in other cases, we expect that the observed data provides important, yet partial, information about scale such that ${\mbox{Id}}_{W^{{\tiny\mbox{tot}}}}(P)\subsetneq\mathcal{W}^{\tiny\mbox{tot}}$. For example, consider the political polling scenario at the start of this article. If the $n$-th poll surveys 10,000 people, we can infer that $W^{{\tiny\mbox{tot}}}_{n}\geq 10,000$. We call these partial compositional surveys.
4.1.3 Relating Definitions back to SRI and Real Problems
In Supplementary Section B.2, we provide a number of theoretical results which establish how estimands and compositional surveys combine to determine if an estimation problem is Full SRI, Partial SRI, or SII. A particularly intuitive result is that, in the setting of a full compositional survey, if an estimand is fully scale reliant, partially scale reliant, or scale invariant, then the resultant estimation problem is Full SRI, Partial SRI, or SII, respectively.
To relate back to real problems addressed in existing literature, recall that calling the data compositional is effectively modeling the observed data as an observation of a full compositional survey. Within the analysis of human microbiota, we find examples of researchers calling these data compositional, yet they are trying to use these data to estimate all three of the estimands presented above (Fernandes et al.,, 2014; Friedman and Alm,, 2012; Silverman et al.,, 2017). In short, this distinction between SII, Partial SRI, and Full SRI are not simply theoretical constructs but relate to real problems encountered in the literature.
5 Limits on Scale Reliant Inference
Having formally defined SRI as a problem of identifiability, we can prove fundamental limits on the types of inference that can be performed using data with arbitrary scale. We frame these limits in terms of common Frequentist and Bayesian inferential criteria. We show that there are strict limits on Frequentist inference but none on Bayesian inference. Still, we show that rigorous Bayesian inference requires a uncommon, specialized type of model called a Bayesian PIM. We use the concept of unacknowledged bias to discuss the practical implication of these limits.
5.1 Limits on Frequentist Inference
The following four theorems follow naturally from the definition of SRI as a problem of identifiability. Similar impossibility proofs can be found in other fields that encounter unidentified models (e.g., Gabrielsen, (1978); Dufour, (1997); Gleser and Hwang, (1987)). Formal proofs of each theorem are provided in Supplementary Section B.3.
Theorem 1 (Consistency).
[…]
Theorem 2 (Unbiasedness).
[…]
Theorem 3 (Calibration).
Suppose there exists $\theta\in\Theta$ and $\lambda\in\Lambda$ such that $Id_{\theta}(P)=\Theta$ (Full SRI). There does not exist a non-trivial [^1] Formally defined in Supplementary Section B.3, used to exclude trivial confidence sets that cover the entire parameter space. confidence set $S(\mathbf{Y})$ calibrated at level $01-\alpha>$ 0.
Theorem 4 (Type-I Error Control).
Suppose there exists some $\theta\in\Theta$ and $\lambda\in\Lambda$ such that $Id_{\theta}(P)=\Theta$ (Full SRI). Then, for any $\alpha>0$, there does not exist a non-trivial [^2] Formally defined in Supplementary Section B.3, used to exclude trivial tests that never reject the null hypothesis. level-$\alpha$ test of the null hypothesis $\theta\in\Theta_{0}\subsetneq\Theta$.
To the best of our knowledge, these are the first formal proofs of what can and cannot be done when analyzing data with arbitrary scale. These theorems provide formal justification for CoDA’s insistence on scale invariance, yet they also show that that existing arguments within the CoDA community are overly restrictive. For example, consistent with CoDA’s arguments, these theorems show that the Frequentist ideal is not attainable for Full SRI problems, a non-trivial result. This contradicts a growing body of literature which claims to have developed rigorous methods for performing DA analysis from compositional data (e.g., Srinivasan et al., (2020); Love et al., (2014); Fernandes et al., (2014)). As we discussed in the prior section, this is a Full SRI problem. Theorem 2 proves that this is not possible. In Supplementary Section C, we review one such paper and discuss how to reconcile those claims with our results. Yet, these theorems also show that rigorous Frequentist inference is still possible in other types of SRI problems, namely Partial SRI problems. In fact, we have recently developed Frequentist methods for Differential Set Analysis (McGovern et al.,, 2023). In short, we have proven that there are some types of scale reliant estimation tasks that can be rigorously performed (relative to Frequentist ideals) using compositional data and demonstrated this claim in McGovern et al., (2023). In the next subsection, we expand beyond Frequentist inference and show that Bayesian inference is possible for all forms of SRI and SII.
5.2 Limits on Bayesian Inference
The previous section established strict limits on Frequentist inference in the setting of SRI. There are no equivalent limits on Bayesian inference. So long as the chosen likelihood satisfies the likelihood principle and the prior is calibrated to prior beliefs, Bayesian inference is possible. Yet, for SRI, satisfying these two criteria can be challenging and requires a specialized type of model called a Bayesian PIM.
For concreteness, consider LFC estimation from a full compositional survey $\mathbf{Y}$ (Full SRI). Suppose our prior beliefs can be represented as a non-singular distribution $p(\theta=(\theta_{1},\dots,\theta_{D}))$. Using Equation 2, this prior can alternatively be written as $p(\theta^{\tiny\mbox{comp}},\theta^{\tiny\mbox{tot}})$ where $\theta^{\tiny\mbox{comp}}$ is a vector with elements $\theta^{\tiny\mbox{comp}}_{d}$ which are functions of $\mathbf{W}^{{\tiny\mbox{comp}}}$ while $\theta^{\tiny\mbox{tot}}$ is a scalar that is a function of $W^{{\tiny\mbox{tot}}}$. To satisfy the likelihood principle, we must choose a likelihood model that accurately reflects the evidence of the observed data $\mathbf{Y}$ relevant to the model parameters $\theta$. Since $\mathbf{Y}$ provides no information about $W^{{\tiny\mbox{tot}}}$, our likelihood must satisfy
$\mathcal{L}(\theta^{\tiny\mbox{comp}},\theta^{\tiny\mbox{tot}};\mathbf{Y})\propto\mathcal{L}(\theta^{\tiny\mbox{comp}};\mathbf{Y}).$
This likelihood, combined with this prior, is an example of a Bayesian PIM: it is a model in which the likelihood, even asymptotically, does not uniquely identify the parameters.
5.3 So What? Unacknowledged Bias
Considering that all models are wrong (Box,, 1976), it is natural to question the practical importance of the above limits: what is so bad about using a Bayesian model that is not a Bayesian PIM? To facilitate this discussion, we borrow the concept of unacknowledged bias from the Bayesian PIM literature (Gustafson,, 2015, pg. 35). Unacknowledged bias is a Frequentist concept which we, like prior literature (e.g., Gustafson, (2015)), choose to apply to both Bayesian and Frequentist methods. Like prior authors, we do not attempt to formally define unacknowledged bias, instead considering it only as a conceptual tool.
Suppose we had chosen a likelihood that was fully identified, e.g.,
$\mathcal{L}(\theta^{\tiny\mbox{comp}},\theta^{\tiny\mbox{tot}};\mathbf{Y})\not\propto\mathcal{L}(\theta^{\tiny\mbox{comp}};\mathbf{Y}).$
Assuming standard regularity conditions (e.g., $p(\theta)>0$ for all $\theta\in\Theta$), the posterior will concentrate to a point mass centered on the unique maximum likelihood estimate asymptotically. Although the model implies complete certainty in the estimate (zero variance), we have already established that the maximum likelihood estimate is incorrect (inconsistent and biased). This combination, an estimator that is biased yet fails to account for the bias through its variance, is called unacknowledged bias (Gustafson,, 2015). A hallmark of unacknowledged bias is a Type-I or Type-II error rate that increases with more data (see Supplemental Figure 2 for visual explanation). It is difficult to do rigorous science when using an estimator with unacknowledged bias. To use such an estimator, we must consider whether we have collected too much data to trust the results of our analysis. In Supplemental Section D, we discuss three common pitfalls encountered when developing Bayesian models for SRI including the use of Poisson or Negative Binomial likelihoods. We provide examples of these pitfalls within the existing literature and discuss how they lead to unacknowledged bias by violating either the likelihood or prior calibration principle.
In SRI, we cannot avoid bias (Theorem 2). Remarkably, Bayesian PIMs can acknowledge their bias and thereby minimize spurious scientific conclusions. Returning to our LFC example, consider that the posterior of the Bayesian PIM will factor as
$p(\theta^{\tiny\mbox{tot}},\theta^{\tiny\mbox{comp}}\mid\mathbf{Y})$ $=p(\theta^{\tiny\mbox{tot}}\mid\theta^{\tiny\mbox{comp}},\mathbf{Y})p(\theta^{\tiny\mbox{comp}}\mid\mathbf{Y})$
$=p(\theta^{\tiny\mbox{tot}}\mid\theta^{\tiny\mbox{comp}})p(\theta^{\tiny\mbox{comp}}\mid\mathbf{Y}).$
In short, the conditional relationship between $\theta^{\tiny\mbox{tot}}$ and $\theta^{\tiny\mbox{comp}}$ is the same before and after having seen the data. As a result, any learning that might occur regarding $\theta^{\tiny\mbox{tot}}$ after observing $\mathbf{Y}$ occurs indirectly through our updated beliefs about $\theta^{\tiny\mbox{comp}}$. So long as the variance of $p(\theta^{\tiny\mbox{tot}}\mid\theta^{\tiny\mbox{comp}})$ is non-zero, the posterior will not concentrate to a point mass asymptotically. Instead, the posterior will acknowledge our lack of information about $\theta^{\tiny\mbox{tot}}$ and retain non-zero variance. This feature of Bayesian PIMs is why they are said to display an acknowledged rather than unacknowledged bias (Gustafson,, 2015). In Section 7, we demonstrate the practical advantages of acknowledging bias in the context of simulated and real data.
6 Scale Simulation Random Variables (SSRVs)
Moving beyond limits on inference, we concentrate on practical solutions to SRI. We focus our attention on Bayesian approaches as they can be applied to both Partial and Full SRI tasks. Two challenges confront the use of Bayesian PIMs for SRI: model specification and computational tractability.
We expect that, in most cases, it will be challenging to specify a likelihood model $\mathcal{L}(\theta;\mathbf{Y})$ that encodes the limited information about $\theta$ provided by $\mathbf{Y}$. For example, in general cases, we expect the estimand will not permit the same type of additive decomposition (i.e., $\theta_{d}=\theta^{\tiny\mbox{comp}}_{d}+\theta^{\tiny\mbox{tot}}$) as it does for LFC estimation (see Supplementary Section B.1 for an example related to covariance estimation). Beyond the likelihood, standard approaches to prior specification in Bayesian PIMs can imply unreasonable assumptions about the system scale. For example, for LFC estimation a prior of the form $p(\theta)$ may make implicit assumptions about $\theta^{\tiny\mbox{tot}}$ though the conditional $p(\theta^{\tiny\mbox{tot}}\mid\theta^{\tiny\mbox{comp}})$. In Section 7.2, we demonstrate how standard priors can lead to elevated rates of Type-I and Type-II errors in practice.
Even after the likelihood and prior are specified, Bayesian PIMs can be challenging to infer. Returning to LFC estimation, consider that the posterior factor $p(\theta^{\tiny\mbox{tot}}\mid\theta^{\tiny\mbox{comp}})$ retains the variance of the prior while the variance of the factor $p(\theta^{\tiny\mbox{comp}}\mid\mathbf{Y})$ will typically decrease with increasing sample size. This leads to substantial anisotropy (i.e., $\mathrm{var}(\theta^{\tiny\mbox{comp}}\mid\mathbf{Y})<<\mathrm{var}(\theta^{\tiny\mbox{tot}}\mid\mathbf{Y})$) which poses computational challenges for traditional Bayesian inferential tools (Gustafson,, 2015). While prior authors have suggested that tools like importance sampling may be used to address these problems, in our experience, those methods work poorly for SRI where posteriors are often high-dimensional.
To address these challenges, we introduce Scale Simulation Random Variables (SSRVs), a modeling framework that can help specify and infer Bayesian PIMs for SRI. We define SSRVs in the following section.
6.1 Defining SSRVs
We start by defining SSRVs before discussing how they relate to Bayesian PIMs and address the aforementioned challenges. For simplicity, we limit our discussion to SSRVs for full compositional surveys. Generalizations to partial compositional surveys are straightforward yet may be more computationally challenging.
We define SSRVs as the combination of three components. The first is the target estimand, $\theta(\mathbf{W}^{{\tiny\mbox{comp}}},W^{{\tiny\mbox{tot}}})$. This estimand also defines the problem as SII versus SRI. The second is the measurement model, $p(\mathbf{W}^{{\tiny\mbox{comp}}}\mid\mathbf{Y})$. The measurement model represents posterior uncertainty in the system composition. The third component is a scale model, $p(W^{{\tiny\mbox{tot}}}\mid\mathbf{W}^{{\tiny\mbox{comp}}})$. The scale model is a prior distribution which represents uncertainty in the system scale and may be conditioned on the system composition. With these components, a SSRV is a random variable $\theta$ with probability distribution $p(\theta\mid\mathbf{Y})$ that can be defined either algorithmically or through an integral transform.
Definition 2 (SSRVs: Algorithm).
The following algorithm simulates $S$ samples from a SSRV defined by the target estimand, measurement model, and scale model:
Data: $\mathbf{Y}$
Result: $S$ samples of $\theta$
for $s$ in $\{1,\dots,S\}$ do
[…]
end for
Algorithm 1 Monte-Carlo Algorithm for Simulating Scale Simulation Random Variables
Definition 3 (SSRVs: Integral Transform).
The following integral transformation defines a SSRV for $\theta$ as the transformation of the scale and measurement models through the target estimand:
[…]
Less formally, SSRVs are simulations of the target estimand that incorporate both uncertainty in the system composition and uncertainty in the system scale. This interpretation is echoed in the SSRV mean:
$\mathbb{E}[\theta\mid\mathbf{Y}]=\mathbb{E}_{p(W^{{\tiny\mbox{tot}}}\mid\mathbf{W}^{{\tiny\mbox{comp}}})}[\mathbb{E}_{p(\mathbf{W}^{{\tiny\mbox{comp}}}\mid\mathbf{Y})}[\theta(\mathbf{W}^{{\tiny\mbox{comp}}},W^{{\tiny\mbox{tot}}})]].$
6.2 Relationship to Bayesian PIMs
SSRVs are Bayesian PIMs of the form $p(\theta\mid\mathbf{Y})$ which are specified using augmented variables $W^{{\tiny\mbox{tot}}}$ and $\mathbf{W}^{{\tiny\mbox{comp}}}$. Consider that, within the integral of Equation 7, the term $p(W^{{\tiny\mbox{tot}}}\mid\mathbf{W}^{{\tiny\mbox{comp}}})p(\mathbf{W}^{{\tiny\mbox{comp}}}\mid\mathbf{Y})$ is the posterior of a Bayesian PIM. Unlike the Bayesian PIMs discussed previously, the Bayesian PIM within the integral of Equation 7 uses $W^{{\tiny\mbox{tot}}}$ and $\mathbf{W}^{{\tiny\mbox{comp}}}$ as parameters rather than $\theta$. SSRVs then use the target estimand to transform this posterior over $\mathbf{W}^{{\tiny\mbox{comp}}}$ and $W^{{\tiny\mbox{tot}}}$ into a posterior over $\theta$. SSRVs are simply an alternative method of specifying Bayesian PIMs for SRI.
SSRVs mitigate the challenges of Bayesian PIMs discussed in Section 6. Using SSRVs, analysts do not need to directly specify a partially identified likelihood $\mathcal{L}(\theta;\mathbf{Y})$. Instead, analysts have the comparably easier task of specifying a fully identified likelihood $\mathcal{L}(\mathbf{W}^{{\tiny\mbox{comp}}};\mathbf{Y})$ when building the measurement model. We expect this will often be a straight-forward compositional modeling task where analysts can make use of a wide variety of available tools (e.g., Silverman et al., (2022)). Beyond the likelihood, the scale model of an SSRV provides a clear picture of modeling assumptions about scale. In Section 7.2 we highlight the benefit of this transparency using real and simulated data examples. SSRVs can also simplify posterior computations due to Algorithm 1 as the measurement model is a standard (identified) Bayesian model, and the scale model is simply a prior. Sampling from both distributions is comparatively straightforward and avoids the anisotropy common with Bayesian PIMs.
Case Study: ALDEx2 is a SSRV with a Degenerate Scale Model
While there are many popular tools used for differential abundance analysis, we highlight ALDEx2 (Fernandes et al.,, 2014) as it has a particularly simple mathematical form which aligns closely with SSRVs. In viewing ALDEx2 as a SSRV, we identify the origin of its unacknowledged bias (demonstrated empirically in later sections) and identify how to improve the model and mitigate that bias.
In Supplementary Section D.3, we present the ALDEx2 model as originally published in Fernandes et al., (2014). For the sake of brevity and clarity, we briefly review a simplified form of the ALDEx2 model which can be described in three steps. In the first step, posterior draws from independent multinomial-Dirichlet models are used to simulate uncertainty in the composition conditioned on the observed data. In the second step, those draws are normalized using the Centered Log-Ratio (CLR) transform which is a transform from the $D$-simplex ($\mathbb{S}^{D}$) to $\mathbb{R}^{D}$ defined as $\phi(x)=\log x-\mathbf{1}\operatorname*{mean}(\log x)$ for a vector $x\in\mathbb{S}^{D}$. In the third step, LFCs are calculated using the CLR-normalized relative abundances. The result is a distribution over LFCs. In fact, these three steps correspond to the measurement model, scale model, and target estimand of a SSRV.
The multinomial-Dirichlet samples imply a measurement model of the form:
$\mathbf{W}^{{\tiny\mbox{comp}}}_{\cdot n}\sim\text{Dirichlet}(\mathbf{Y}_{\cdot n}+\alpha)$
where $\alpha$ are the chosen Dirichlet prior parameters (default to $\alpha=0.5\cdot\mathbf{1}_{D}$ in ALDEx2). The target estimand is the LFC estimand (Eq. 1):
$\theta_{d}=\operatorname*{mean}_{n:x_{n}=1}\log\mathbf{W}_{dn}-\operatorname*{mean}_{n:x_{n}=0}\log\mathbf{W}_{dn}.$
Together, these two components suggest that the CLR normalization step represents some form of implicit scale model. This normalization can be written as:
$\log\mathbf{W}_{dn}=\log\mathbf{W}^{{\tiny\mbox{comp}}}_{dn}-\operatorname*{mean}\left(\log\mathbf{W}^{{\tiny\mbox{comp}}}_{\cdot n}\right).$
Comparing this to $\log\mathbf{W}_{dn}=\log\mathbf{W}^{{\tiny\mbox{comp}}}_{dn}+\log W^{{\tiny\mbox{tot}}}_{n}$ implied by Equation 2, it is clear that this normalization equates to a scale model of the form:
$\log W^{{\tiny\mbox{tot}}}_{n}\sim\delta[-\operatorname*{mean}\left(\log\mathbf{W}^{{\tiny\mbox{comp}}}_{\cdot n}\right)]$
where $\delta[x]$ denotes a point mass centered at the point $x$. ALDEx2 does not account for any uncertainty in its implicit assumptions regarding the unmeasured system scale. Errors in this assumption lead to biased estimates.
Beyond sequence counting, we expect that unacknowledged bias due to errors in implicit assumptions regarding scale may be found in other fields. In Supplementary Section F, we provide an example of an implicit assumption in a specific type of ecological survey.
7 Empirical Studies of Sequence Counting
Through the analysis of two case-studies, we show that popular tools for differential abundance analysis display Type-I and/or Type-II error rates due to unacknowledged bias originating from problems of scale. Building on our description of ALDEx2 as a SSRV, we show that these error rates can be drastically reduced by considering error in assumptions about scale (i.e., using a scale model with non-zero variance).
7.1 ALDEx2’s Degenerate Scale Model Drives False Positives
We simulated a microbiome study comparing human gut microbiota in otherwise healthy individuals before and after treatment with a mild, narrow spectrum antibiotic whose true effect is to slightly decrease the abundance of 4 of the 20 simulated taxa (full simulation details provided in Supplementary Section G.1). Figure 1a shows the simulated true abundances ($\mathbf{W}$) for a set of $N=100$ study participants equally split between the pre- and post-antibiotic conditions. We simulated data $\mathbf{Y}$ with arbitrary scale by resampling the true abundances $\mathbf{W}$ with a multinomial distribution so the arbitrary scale of each sequenced sample $\mathbf{Y}_{\cdot n}$ has a total of $5,000$ total counts. We applied two popular DA models (DESeq2 (Love et al.,, 2014) and ALDEx2 (Fernandes et al.,, 2014)) to $\mathbf{Y}$ and, in Figure 1, show the pattern of true/false positives/negatives that result. In terms of Type-I Error control, both methods are worse than determining statistical significance by flipping a fair coin. Moreover, in Supplementary Figure 3 we show that the false discovery rate of both methods increases with increasing sample size – the hallmark of unacknowledged bias.
Figure 1: (a) Modeling results for DESeq2, ALDEx2, and a SSRV model. The Relaxed scale model is centered on the same geometric assumption as CLR normalization assumption but considers error in that assumption by using $\alpha=0.5$. b) Sensitivity analysis showing the estimated standardized log-fold change (LFC) for each taxa under the Relaxed model as a function of $\alpha$. The average and standard deviation of log fold change was computed across Monte Carlo samples, and a standardized value was computed for each entity by dividing the mean LFC by its corresponding standard deviation. A black line signifies that the estimated effect size was statistically different from zero at 95% confidence. The four taxa that were truly differentially expressed are labeled (Taxa 3, 4, 15, and 20).
As described in the prior section, we expect that the implicit scale model used by ALDEx2 underestimates a researcher’s true uncertainty in the system scale leading to an unacknowledged bias and driving the false positives seen in Figure 1. To evaluate this hypothesis, we created a relaxation of the ALDEx2 model replacing the degenerate (point-mass) scale model with the following relaxation: $\log W^{{\tiny\mbox{tot}}}_{n}\sim N(-\operatorname*{mean}\left(\log\mathbf{W}^{{\tiny\mbox{comp}}}_{\cdot n}\right),\alpha^{2})$.
Our relaxation is sufficient to remove all of ALDEx2’s false positives and retain most if not all of the true positives (Figure 1a). Additionally, our relaxation addresses the unacknowledged bias and does not lead to the same asymptotic increase in false discovery rates as seen with ALDEx2 and DESeq2 (Supplemental Figure 3). Moreover, sensitivity analysis shows that our results are largely insensitive to the choice of $\alpha$: even small amounts of uncertainty are sufficient to remove the majority of false positives at the cost of just a single false negative (taxa 3; Figure 1b). Only when $\alpha$ is chosen to be implausibly large ($\alpha>1.4$) does the relaxed ALDEx2 model fail to find true positives. Considering that the CLR estimate of scale corresponds to $\hat{\theta}^{\perp}_{CLR}=0.03$ a value of $\alpha>1.4$ reflects a greater than 5% probability that the log-fold-change of scales, $\theta^{\perp}$, is outside of the range $[\exp^{-1.96\times 1.4^{2}+\hat{\theta}^{\perp}_{CLR}},\exp^{1.96\times 1.4^{2}+\hat{\theta}^{\perp}_{CLR}}]=[0.02,48.02]$; which is implausible for a narrow spectrum antibiotic. In summary, our results indicate that the CLR normalization in ALDEx2 drives its observed false positives and that these can be mitigated by considering potential errors in assumptions about scale. The observed pattern of false positives within DESeq2 suggests similar effects may be at play within that model even though it uses a different median-of-ratios normalization.
7.2 Application to a novel 16S rRNA-Seq dataset
Although artificial gut systems (bioreactors) are seeded with human feces and often used as a model for host-associated microbiota, many bioreactors undergo a two week equilibration period in which the inoculated communities diverge from those seen in human hosts and adapt to their new ex vivo environment. To evaluate the limits of these model systems, we ran and collected data from 8 bioreactors over the course of the first two weeks after inoculation of which 6 were used for analyses due to experimental constraints (see Supplemental Section G.2). Each bioreactor was seeded with 100mL of fecal slurry and maintained at a constant 400mL of growth media over the course of the experiment. Our primary interest was in determining which of the microbial taxa present in the innoculum were able to grow in the bioreactors. We quantify this interest using the differential abundance target estimator given in Equation (1) with $x_{n}=0$ denoting day 1 of the experiment and $x_{n}=1$ denoting day 14. In this case, we were primarily interested in the de facto total scale, i.e., we wanted to identify taxa that increased or decreased in total cellular number over the two week period.
This data is ideal for our study as our experimental system allowed us to obtain high-quality measurements of scale in this bioreactors. Because each vessel was homogeneously mixed, microbial concentration measured from a sample of the vessel via flow cytometry was presumably identical to microbial concentration within the vessel. Moreover, because each vessel was maintained at a constant 400mL volume throughout the course of the experiment, changes in microbial concentration between vessels or between time-points were proportional to changes in total microbial load. Flow cytometry measurements were obtained to measure total cellular concentration. These measurements were taken in triplicate for each 16S rRNA sequencing measurement to allow us to quantify measurement noise. Further experimental details are presented in Supplemental Section G.2.
Five models were compared on the resultant data: one Bayesian PIM, two SSRVs, ALDEx2, and DESeq2. We used the following multinomial log-normal Bayesian PIM, which is an extension of the Bayesian multinomial logistic-normal models that have become popular for analysis of microbiota survey data (Silverman et al.,, 2022):
$\mathbf{Y}_{\cdot n}$ $\sim\text{Multinomial}(\mathbf{W}^{{\tiny\mbox{comp}}}_{\cdot n})$ (8)
$\mathbf{W}_{\cdot n}$ $=\exp({\psi_{\cdot n}})$ (9)
$\psi_{\cdot n}$ $\sim N(\mathbf{B}X_{\cdot n},\boldsymbol{\Omega})$ (10)
with priors $\mathbf{B}\sim N(\mathbf{M},\boldsymbol{\Omega},\boldsymbol{\Gamma})$ and $\boldsymbol{\Omega}\sim IW(u,\boldsymbol{\Xi})$. This is a Bayesian PIM since the multinomial likelihood is only parameterized in terms of $\mathbf{W}^{{\tiny\mbox{comp}}}$ whereas prior is defined over both $\mathbf{W}^{{\tiny\mbox{comp}}}$ and $W^{{\tiny\mbox{tot}}}$. We chose the prior hyperparameters to echo standard defaults we have observed in the literature, i.e., we chose $\mathbf{M}=0$ and choose $\mathbf{\Gamma}$ and $\mathbf{\Xi}$ to be diagonal matricies (See Supplementary Section G.2 for prior specifications).
There are two critical limitations of this PIM. First, it is computationally infeasible to infer this model directly. Instead, to obtain posterior samples of the parameter $\mathbf{B}$, we had to derive an equivalent SSRV and sample $\mathbf{B}$ using Algorithm 1. Supplementary Section H derives the SSRV representation of this Bayesian PIM and shows that the measurement model is a Bayesian multinomial logistic-normal model and the scale model is a log-normal linear model. The second, more important, limitation of this model is that the assumptions about scale are not transparent. In fact, this seemingly innocuous and standard prior specification actually implies strong, biologically implausible assumptions: implicit in this prior is a non-trivial conditional relationship of the form $p(\theta^{\tiny\mbox{tot}}\mid\theta^{\tiny\mbox{comp}})$. This leads to a posterior for $\theta^{\tiny\mbox{tot}}$ which reflects a strong, biologically unreasonable belief that total microbial load is higher on day 1 than on day 14 (see Supplementary Section G.2). Below we show this leads to unacknowledged bias which presents as elevated Type-I and Type-II error rates.
To illustrate how SSRVs can enhance model specification and ease prior specification, we designed two SSRVs using the same measurement model as the above Bayesian PIM but with two different scale models. Letting $n\in\{1,\dots,N\}$ denote a given vessel at a given time-point, these two scale models were both log-normal models of the form :
$\log W^{{\tiny\mbox{tot}}}_{n}\sim N(c+\mu_{n},\alpha_{n}^{2})$ (11)
where $c$ is some constant and $\mu$ is subject to the constraint $\sum_{n=1}^{N}\mu_{n}=0$. Note that the LFC target estimand given by Equation 1 is invariant to the value of $c$ and so we can set $c=0$ with no loss of generality.
We created the Design scale model based only on prior knowledge of the experimental design. Since we know that each vessel contained 400mL of growth media and the innoculum consisted of 100 mL of fecal slurry, we expect that the total microbial load would, on average, increase from the day 1 to day 14 by a factor of approximately 4. We encapsulated these beliefs in the scale model
$\frac{f(W^{{\tiny\mbox{tot}}}_{{j},\text{t=1}})}{f(W^{{\tiny\mbox{tot}}}_{{j},\text{t=14}})}\sim\mathrm{LogNormal}(\bar{d},\tau^{2})$
where $\bar{d}=\log(100/400)$ and $\tau=0.5$.
Finally, we created the Flow scale model built from the flow-cytometry measurements. In this model $\mu_{n}$ and $\alpha^{2}_{n}$ were estimated as mean and variance our flow cytometry measurements on the log scale. We consider the SSRV created using this scale model our gold standard against which we define false positives. For concreteness, we define a result as a false positive in two cases: first, that the result is significant in a given model but not in the gold standard and second, that the result is significant in both the model and the gold standard, but the two models disagree on sign.
We applied the Bayesian PIM, the two SSRVs, ALDEx2, and DESeq2 to the artificial gut data; the results are summarized in Figure 2. As before we find that ALDEx2 and DESeq2 have high inferential error rates. However, unlike the prior study where these were primarily Type-I errors, here we observe far more Type-II errors owing to the fact that most taxa are actually increasing in abundance as the community grows to occupy the larger vessels and this study has a relatively small sample size ($N=12$). Still we find the hallmark of unacknowledged bias: using resampling, Supplemental Figure 5 shows that ALDEx2 and DESeq2 display increasing rates of false discoveries with increasing sample size while consistently displaying Type-II error rates far in excess of the Design model. For example with just 20 vessels, the FDR of ALDEx2 and DESeq2 are approximately 30% and 15% respectively. In contrast, even at 50 vessels, the Design model controls FDR at a nominal 5%. As expected based on our investigation of the Bayesian PIM prior, this model displays a false discovery 45% which increases with more data (Supplemental Figure 5).
Finally, Figure 3 shows that the major conclusions of this analysis, namely the remarkably performance improvement of the Design model compared to alternatives is largely insensitive to the choice of the parameter $\tau$. So long as $\tau$ is greater than 0.35, the FDR of the Design model will be less than 3.5%. Moreover, so long as we choose $\tau$ less than 5, the Type-II error rate of the design-based SSRV will be less than that of either DESeq2 or ALDEx2. A value of $\tau\geq 5$ would require we accept, with 2.5% probability, that the microbial load at day 14 was at least $4\exp(1.96*5^{2})\approx 7.6\times 10^{21}$ times greater than at day 1. Even if there was only one microbe at Day 1, this is not physically possible: as bacterium weights approximately $10^{-12}$ grams, this would require the vessels at day 14 to weight more than $10^{6}$ kilograms.
Figure 2: Comparison of results from ALDEx2, DESeq2, the Bayesian PIM, and two SSRVs using real data from an ex vivo artificial gut study. “GS” denotes the gold-standard model against which true/false positives/negatives are defined. Our goal standard (the SSRV using the Flow scale model) is the only model that could use the replicate flow-cytometry measurements collected in this study. For brevity, taxa that were identified as differentially abundant with respect to any of the 5 models are not shown (based on p-value less than 0.05 or posterior 95% credible interval not covering zero).
Figure 3: Sensitivity analysis of the Design SSRV as a function of the parameter $\tau$ in the Design scale model. The mean, standard deviation, and 95% credible interval were computed using the posterior distribution of log-fold change for each entity. The standardized log-fold change (LFC) was calculated by dividing the mean log-fold change by the corresponding estimate for standard deviation. The standardized LFC as a function of $\tau$ is shown for each taxa in the artificial gut study and is colored black if the corresponding credible interval excluded zero and grey otherwise. As an illustration, we highlight the 5 taxa that are the most clearly differentially abundant in the sense that their credible sets exclude zero over the largest range of $\tau$. For comparison, we depict the value of $\tau$ used in Figure 2 ($\tau=0.5$) as a black dotted line.
8 Discussion
In this article, we have framed the challenge of data with arbitrary scale as an estimation problem involving partially identified models. Through this formulation, we have been able to provide the first formal limits on inference with these data. Our results are consistent with the CoDA perspective: they show that care is needed as spurious scientific conclusions can be reached due to the lack of information about the system scale contained in these data. However, unlike the CoDA perspective, our results argue that a dogmatic restriction to scale invariance is overly prohibitive. Consider that, even if data is compositional (i.e., a full compositional survey with simplex valued measurements), Bayesian PIMs can be used to perform rigorous SRI. In fact, even Frequentist hypothesis testing and confidence estimation can be done rigorously when analyzing compositional data if the estimand is partially scale reliant. Overall, we argue that SRI is possible yet requires new types of statistical methods that can account for the limited information contained in the observed data.
Beyond discussing limits on inference, this article also introduces the class of Bayesian PIMs for SRI. There is an interesting connection between the SSRV formulation of Bayesian PIMs and adversarial robustness. Adversarial methods are well known in the machine learning literature as a way of enhancing the robustness or stability (Yu,, 2020) of predictions. For example, within computer vision, noise is sometimes added to images in order to ensure that trained models are stable and are not sensitive to slight variation when applied out-of-sample (Santos and Papa,, 2022). Due to concerns about rigor and reproducibility in science, some authors have advocated that adversarial perturbations can be applied to ensure robustness of statistical inference (Yu,, 2020). Within SSRVs, the use of the scale model can be viewed as a type of adversarial perturbation: for each sample from the SSRV we apply a different random perturbation sampled from the scale model. One way of understanding the utility of Bayesian PIMs (and SSRVs) is that they involve a form of adversarial robustness to perturbations of scale.
In this article, we have focused on Bayesian approaches to SRI as these methods, unlike Frequentist methods, remain valid even in the setting of Full SRI. To clarify the problem with Frequentist methods, consider the problem of confidence estimation. Intuitively, a level-$\alpha$ confidence set for an estimand $\theta$ should, with probability $1-\alpha$ cover the identification region ${\mbox{Id}}_{\theta}(P)$. The problem is that, for Full SRI, ${\mbox{Id}}_{\theta}(P)=\Theta$ and therefore any level-$\alpha$ confidence set must be uselessly large. Yet, for Partial SRI, useful confidence estimation (and hypothesis testing) may be possible as ${\mbox{Id}}_{\theta}(P)\subsetneq\Theta$. For example, we recently proved that useful and robust hypothesis testing is possible for a particular partial SRI problem called differential set analysis (McGovern et al.,, 2023). Overall, we believe that Frequentist SRI for Partial SRI problems is an intriguing and likely fruitful area for future research. In fact, there is already a rich literature on Frequentist confidence estimation for partially identified models can likely be adapted to the task. We refer interested readers to Imbens and Manski, (2004) and Stoye, (2009) for a thorough review of those methods.
For simplicity, we have focused on compositional surveys which are a subset of data with arbitrary scale where the system composition is identifiable. Yet in many practical applications, we expect that the system composition may not be identifiable as the measurement process may involve unobserved compositional biases. For example, in the 16S rRNA microbiome surveys, the PCR amplification process can lead to an over-representation of some taxa compared to others in the sequenced pool (Silverman et al.,, 2021). If compositional bias is substantial, then it may tighten the limits on inference further restricting the types of inference that can be performed rigorously with these data. Just as Bayesian PIMs can be used to address arbitrary scale, we expect these same models may be adapted to provide robustness to compositional bias. Overall, we believe that the study of compositional bias represents a fruitful area for future research.
Acknowledgments and Funding Sources
We thank Rachel Silverman, Juan Jose Egozcue, Vera Pawlosky-Glahn, Gregory Gloor, Michael Love, and James Morton for their manuscript comments. JDS and MPN were supported in part by NIH 1R01GM148972-01. LAD acknowledges support from the the Global Probiotics Council, a Searle Scholars Award, the Hartwell Foundation, an Alfred P. Sloan Research Fellowship, the Translational Research Institute through Cooperative Agreement NNX16AO69A, the Damon Runyon Cancer Research Foundation, and NIH 1R01DK116187-01. SM would like to acknowledge the support of grants NSF IIS-1546331, NSF DMS-1418261, NSF IIS-1320357, NSF DMS-1045153, and NSF DMS-1613261.
Data Access and Code Availability
All data and R scripts necessary to replicate all analyses in the manuscript and supplement are located at https://github.com/michellepistner/ScaleSimulation.