### v5

# Scale Reliant Inference

Abstract
Many scientific fields, including human gut microbiome science, collect multivariate count data where the sum of the counts is unrelated to the scale of the underlying system being measured (e.g., total microbial load in a subject’s colon). This disconnect complicates downstream analyses such as differential analysis in case-control studies. This article is motivated by a novel study of in vitro human gut microbiome models. Popular tools for analyzing these data led to dramatically elevated rates of both false positives and false negatives. To understand those failures, we provide a formal problem statement that frames these challenges of scale in terms of the classical theory of identifiability. We call this the problem of Scale Reliant Inference (SRI). We use this formulation to prove fundamental limits on SRI in terms of criteria such as consistency and type-I error control. We show that the failures of existing methods stem from a fundamental failure to properly quantify uncertainty in the system scale. We demonstrate that a particular type of Bayesian model called a Bayesian Partially Identified Model (PIMs) can correctly quantify uncertainty in SRI. We introduce Scale Simulation Random Variables (SSRVs) as a flexible and efficient approach to specifying and inferring Bayesian PIMs. In the context of both real and simulated data, we find SSRVs drastically decrease type-I and type-II error rates.
Multivariate Count Data ,
Partially Identified Models,
Compositional Data Analysis,
Microbiome Data,
keywords:
[…]
, , , , , , and
1 Introduction
This article is motivated by statistical challenges that arise in many areas of science and have become a central focus of modern microbiome research. Microbiome sequence count data are typically represented as an $D\times N$ table of counts $\mathbf{Y}$ with elements $\mathbf{Y}_{dn}$ representing the number of DNA sequences observed that map to the $d$-th microbial taxon in the $n$-th biological sample (e.g., a fecal sample from the $n$-th study participant). The measurement process is non-trivial yet, for the purposes of the present article, can be thought of as a multivariate sample: there is a large pool of microbes in the human gut that could be measured, and our measurements capture a small, arbitrary-sized sample of that pool (Silverman et al., 2018; Roche and Mukherjee, 2022). By arbitrary sized, we allude to the fact that the size of the sample (the sequencing depth $\sum_{d}\mathbf{Y}_{dn}$) is typically independent of the scale of the biological system from which that sample was obtained (e.g., the total number of microbes in the subject’s colon) (Lovell et al., 2011; Props et al., 2017; Vandeputte et al., 2017). As a result, many authors call these data compositional, alluding to the idea that the data only provides information about the proportional abundance of each microbe and not their absolute abundance (Gloor et al., 2016a; Lovell et al., 2011; Gloor et al., 2017). The problem is that many of the scientific questions asked of these data require information about absolute abundances rather than just proportional abundances.
For concreteness, this article introduces a novel microbiome study that aimed to quantify the limits of in vitro artificial gut vessels as models of human gut microbiota. Eight artificial gut systems were inoculated with a human fecal microbial community sample. The inoculated vessels incubated those communities for two weeks. Our goal was to identify which of the $D$ microbial taxa in the community can grow within the vessels, quantified as an increased abundance on day 14 compared to day 1. This is an example of a differential abundance analysis which aims to identify taxa that differ in abundance between two experimental conditions. For clarity, we are not interested in changes in proportional abundances as, for example, a taxon may decrease in relative abundance between days 1 and 14 simply because its growth was outpaced by some other microbe that grew more rapidly. While differential abundance analyses are performed in many, if not most, microbiome studies (Lin and Peddada, 2020), we find that popular tools for the task lead to exceedingly high rates of false positives and false negatives in our study. In this article, we argue that this is not simply an issue of data size, measurement error, or choice of parametric model but a fundamental failure of these methods to properly quantify uncertainty originating from the lack of scale information in these data.
The main contributions of this article are the following. First, we provide a formal problem statement that defines what it means to ask a question that depends on scale using data that lacks information about scale. We call this the problem of Scale Reliant Inference (SRI) and frame it in terms of the classical theory of identifiability. We discuss the advantages of our formulation over the related problem of Compositional Data Analysis (CoDA). Second, we use this formulation to prove fundamental limits on SRI. Contrary to numerous published claims, we prove that classical inferential criteria such as consistency and Type-I error control cannot be attained in many SRI problems. In contrast, we show that Bayesian inferential criteria can be achieved by a specialized class of models called Bayesian Partially Identified Models (PIMs). We introduce a simulation framework called Scale Simulation Random Variables (SSRVs), which can simplify the specification and inference of Bayesian PIMs for SRI. Finally, we demonstrate SSRVs through both simulated and real data analysis; in particular, we use SSRVs to tackle the differential abundance analysis at the core of the artificial gut vessel study introduced above. In the context of real data, we find SSRVs control false discovery rates while popular alternatives display rates as high as 30 percent. Moreover, SSRVs drastically reduce false rates, identifying over three times more true positives than alternatives. Our methods have already been implemented into popular software tools (Nixon, Gloor and Silverman, 2024). Moreover, Our findings, namely the dramatic decreases in Type-I and/or Type-II error rates compared to existing tools, has been validated in the context of other microbiome datasets (Nixon, Gloor and Silverman, 2024) and meta-transcriptomics (Dos Santos et al., 2024).
While this article focuses on applications to microbiome analysis, the problems discussed and the theory and methods developed are more general. Beyond microbiome studies, scale reliant inference poses problems in other areas of biomedicine (e.g., gene expression studies) (Fernandes et al., 2014; Lovell et al., 2011), macro-ecology (Billheimer, Guttorp and Fagan, 2001; Jackson, 1997; Murphy et al., 2016), political poling (Topîrceanu and Precup, 2020), demography (Rossi et al., 1987), and epidemiology (Silverman, Hupert and Washburne, 2020). In all these fields, multivariate count data is collected where the total counts are arbitrary and unrelated to the measured systems’ underlying scale (i.e., size).
2 A Theory for Scale Reliant Inference
In this section we introduce a theory for scale reliant inference from data that lacks information about scale. We also discuss implications of this theory on inference procedures. First we set notation that will be used throughout the paper.
2.1 Notation
We use $P_{a,b}$ to denote a probability model which is a function mapping from a parameter space $\mathcal{A}\times\mathcal{B}$ to a set of distributions $\mathcal{P}$. For a random variable $x$ the notation $P_{a,b}(x)$ denotes a specific distribution of $x$ within the set $\mathcal{P}$, a point within the image of the model $P_{a,b}$. The identification region ${\mbox{Id}}_{a,b}(P)$ is the pre-image of the model at a specific distribution $P\in\mathcal{P}$:
${\mbox{Id}}_{a,b}(P)=\{(a,b)\in\mathcal{A}\times\mathcal{B}:P_{a,b}=P\}.$
The model is identified at $P$ if ${\mbox{Id}}_{a,b}(P)$ is a singleton set. We define the marginal identification region of $a$ as
${\mbox{Id}}_{a}(P)=\{a\in\mathcal{A}:\exists\>b\in\mathcal{B},P_{a,b}=P\}.$
The marginal identification region of $b$ (${\mbox{Id}}_{b}(P)$) is defined analogously. Statements about the identifiability of a single parameter within a multi-parameter model refer to the marginal identification region: i.e., $a$ is identified at $P$ if ${\mbox{Id}}_{a}(P)$ is a singleton set. While identification can depend on the specific point $P\in\mathcal{P}$, for brevity, we will often omit this detail when the specific point can be inferred from context.
We consider a $D\times N$ matrix of observed data $\mathbf{Y}$ which is a measurement of a $D\times N$ dimensional underlying system $\mathbf{W}$. We place no mathematical restrictions on $\mathbf{Y}$ and only require that $\mathbf{W}$ is positive-valued (integer or continuous) such that the elements $\mathbf{W}_{dn}$ represent the amount of class $d$ in source $n$ (e.g., the abundance of taxon $d$ in subject $n$’s gut). The system $\mathbf{W}$ can be uniquely decomposed into its composition ($\mathbf{W}^{{\tiny\mbox{comp}}}$, proportions) and scale ($W^{{\tiny\mbox{tot}}}$, totals) via:
$\mathbf{W}_{dn}$ $=\mathbf{W}^{{\tiny\mbox{comp}}}_{dn}\,\,W^{{\tiny\mbox{tot}}}_{n}$ (1)
$W^{{\tiny\mbox{tot}}}_{n}$ $=\sum_{d=1}^{D}\mathbf{W}_{dn}.$
This implies that $\mathbf{W}^{{\tiny\mbox{comp}}}$ is a $D\times N$ matrix with columns that sum to $1$, and $W^{{\tiny\mbox{tot}}}$ is a positive-valued $N$-vector. We denote rows of a matrix with a placeholder dot, e.g., $\mathbf{W}_{\cdot n}$ is the $n$th row of $\mathbf{W}$.
2.2 Formal Problem Statement
This subsection presents a formal definition of Scale Reliant Inference (SRI) followed by an example of SRI in microbiome data analyses.
Assume the observed data $\mathbf{Y}$ are generated following a probability model $P_{\mathbf{W}^{{\tiny\mbox{comp}}},W^{{\tiny\mbox{tot}}}}$. We say that $\mathbf{Y}$ lacks information about scale if the parameter $W^{{\tiny\mbox{tot}}}$ is unidentified at $P_{\mathbf{W}^{{\tiny\mbox{comp}}},W^{{\tiny\mbox{tot}}}}(\mathbf{Y})$. For simplicity, this article assumes that $\mathbf{W}^{{\tiny\mbox{comp}}}$ is identified. We make this assumption to focus our attention on the impact of scale and not other sources of measurement limitations associated with $\mathbf{W}^{{\tiny\mbox{comp}}}$ (such as PCR bias in microbiome studies Silverman et al. (2021)). We denote data generation models where $\mathbf{W}^{{\tiny\mbox{comp}}}$ can be identified but $W^{{\tiny\mbox{tot}}}$ cannot as compositional surveys. We choose this name to highlight the connection to compositional data. That connection, along with important distinctions between SRI and compositional data analysis are reviewed in Section 2.3.
SRI can be framed as learning the map $\mathbf{Y}\mapsto\theta$ where $\theta:=\theta\left(\mathbf{W}^{{\tiny\mbox{comp}}},W^{{\tiny\mbox{tot}}}\right)\in\Theta$. We call $\theta$ a target estimand because it is the quantity one wants to estimate. We also introduce a nuisance function $\lambda=\lambda(\mathbf{W}^{{\tiny\mbox{comp}}},W^{{\tiny\mbox{tot}}})$ to ensure that the following concatenated function $f$ defined as
$(\theta,\lambda)=f(\mathbf{W}^{{\tiny\mbox{comp}}},W^{{\tiny\mbox{tot}}})=(\theta(\mathbf{W}^{{\tiny\mbox{comp}}},W^{{\tiny\mbox{tot}}}),\lambda(\mathbf{W}^{{\tiny\mbox{comp}}},W^{{\tiny\mbox{tot}}}))$
is invertible. With this condition, the model $P_{\mathbf{W}^{{\tiny\mbox{comp}}},W^{{\tiny\mbox{tot}}}}$ can be reparameterized in terms of $(\theta,\lambda)$ such that $P_{\mathbf{W}^{{\tiny\mbox{comp}}},W^{{\tiny\mbox{tot}}}}=P_{f^{-1}(\theta,\lambda)}$. We define SRI in terms of the identifiability of $\theta$ within this reparameterized model.
Definition 2.1 (Scale Reliant and Scale Invariant Inference).
If $\mathbf{Y}$ is a compositional survey, the problem of estimating $\theta$ is Scale Reliant Inference (SRI) if $\theta$ is unidentified at $P_{\theta,\lambda}(\mathbf{Y})$. Otherwise (if $\theta$ is identified), the estimation problem is Scale Invariant Inference (SII).
To help motivate this definition, consider that, if $\mathbf{Y}$ is a compositional survey, the only way that $\theta$ can be unidentified is if it depends on $W^{{\tiny\mbox{tot}}}$ (since $\mathbf{Y}$ provides information about $\mathbf{W}^{{\tiny\mbox{comp}}}$ and $\theta$ is a known function). Therefore, our problem statement provides focus to our study: it restricts attention to the setting where the lack of information about $W^{{\tiny\mbox{tot}}}$ contained in $\mathbf{Y}$ impedes our ability to uniquely identify $\theta$.
Example (Differential Abundance Analysis from Microbiome Data).
The analysis of human gut microbiome data often centers around a $D\times N$ sequence count table $\mathbf{Y}$ with elements $\mathbf{Y}_{dn}$ denoting the number of sequenced (measured) DNA molecules that map to a particular microbial taxon $d$ in fecal sample $n$. For concreteness, we provide a toy dataset ($D=4$) in Supplementary Section A. The measurement process is complex and requires many experimental and computational steps to transform fecal samples into a count table $\mathbf{Y}$ (Caporaso et al., 2011). We model the measurement process as random, there exist some true system with composition $\mathbf{W}^{{\tiny\mbox{comp}}}$ and scale $W^{{\tiny\mbox{tot}}}$ but it is only observed through the noisy measurements $\mathbf{Y}$. While the true number of microbes in a system is integer valued; for simplicity, we assume that those counts are large enough that we can approximate $W_{dn}\in\mathbb{R^{+}}$.
Many authors call these data compositional owing to the measurement process. The measurement process can be thought of as a type of poll: there is a pool of microbes that could be measured, but only an arbitrarily sized subset of those are (Fernandes et al., 2014; Lovell et al., 2011). Moreover, due to complexities such as DNA amplification, it is not even possible to assume that the system has more sequences than the sample ($\mathbf{W}_{dn}>\mathbf{Y}_{dn}$). Overall, the scale of the underlying system is lost through this process leading to data that provides information about the system composition but not the system scale (Vandeputte et al., 2017; Props et al., 2017). We model sequence count data as a compositional survey to reflect this feature of the data.
The goal of many microbiome studies is to determine which, if any, taxa differ in abundance between two mutually exclusive conditions (e.g., disease versus health). Letting $x_{n}\in\{0,1\}$ denote the condition of sample $n$, these differential abundance analyses, often center on estimation of Log-Fold-Changes (LFCs):
$\theta_{d}=\operatorname*{mean}_{n:x_{n}=1}\log\mathbf{W}_{dn}-\operatorname*{mean}_{n:x_{n}=0}\log\mathbf{W}_{dn}.$ (2)
LFCs are therefore the target estimand in differential abundance analyses. Combining Equations (1) and (2) leads to the following relationship
$\theta_{d}$ $=\underbrace{\operatorname*{mean}_{n:x_{n}=1}(\log\mathbf{W}^{{\tiny\mbox{comp}}}_{dn})-\operatorname*{mean}_{n:x_{n}=0}(\log\mathbf{W}^{{\tiny\mbox{comp}}}_{dn})}_{\theta^{\tiny\mbox{comp}}_{d}}+\underbrace{\operatorname*{mean}_{n:x_{n}=1}(\log W^{{\tiny\mbox{tot}}}_{n})-\operatorname*{mean}_{n:x_{n}=0}(\log W^{{\tiny\mbox{tot}}}_{n})}_{\theta^{\tiny\mbox{tot}}}$
$=\theta^{\tiny\mbox{comp}}_{d}+\theta^{\tiny\mbox{tot}}.$ (3)
In words, the LFC estimand requires knowledge of how the proportional abundance of the $d$-th taxon changes between conditions ($\theta^{\tiny\mbox{comp}}_{d}$) as well as how total microbial load changes between conditions ($\theta^{\tiny\mbox{tot}}$). Because microbiome data ($\mathbf{Y}$) is a compositional survey, it follows that $\theta^{\tiny\mbox{tot}}$, and therefore $\theta_{d}$, are not identifiable. By definition, differential abundance analysis from microbiome data is a problem of Scale Reliant Inference.
Before progressing, we discuss a subtlety of our model that is a source of frequent confusion. There are often two sources of uncertainty in SRI. Using LFC estimation from microbiome data as an example, there will will be uncertainty in $W^{{\tiny\mbox{tot}}}$, and therefore $\theta^{\tiny\mbox{tot}}$, since both are unidentified. Yet we will also have uncertainty in $\mathbf{W}^{{\tiny\mbox{comp}}}$, and therefore $\theta^{\tiny\mbox{comp}}$, since we only observe a noisy, finite sample $\mathbf{Y}$ and not $\mathbf{W}^{{\tiny\mbox{comp}}}$ directly. To clarify, our assumption that $\mathbf{W}^{{\tiny\mbox{comp}}}$ is identifiable does not imply that $\mathbf{W}^{{\tiny\mbox{comp}}}$ is known. Identifiability is a feature of the distribution of observables $P_{W^{{\tiny\mbox{tot}}},\mathbf{W}^{{\tiny\mbox{comp}}}}(\mathbf{Y})$; yet, we cannot know that distribution without potentially infinite number of replicate studies. In microbiome data analyses, we can have substantial uncertainty in $\mathbf{W}^{{\tiny\mbox{comp}}}$ due to the variability in the count data $\mathbf{Y}$. Still, this article places comparatively less emphasis on this later source of compositional uncertainty as there is already substantial literature on that subject which we will use in later sections (e.g., Gloor et al. (2016b); Silverman et al. (2018)).
2.3 Comparing Compositional Data Analysis and SRI
This article does not require a detailed understanding of Compositional Data Analysis (CoDA). Still, CoDA has become increasingly popular for analyzing human gut microbiome data, and as a result, we provide a brief comparison for readers who are already familiar with that literature. A broader review of CoDA and a more detailed introduction to CoDA in the context of microbiome data can be found in Greenacre et al. (2023) and Gloor et al. (2016a), respectively.
In brief, CoDA studies simplex-valued data: multivariate, continuous, positive-valued vectors that sum to a constant value. In the context of microbiome data, researchers often identify each datum $\mathbf{Y}_{\cdot n}$ with a single vector on simplex via normalization (e.g., $\mathbf{W}^{{\tiny\mbox{comp}}}_{\cdot n}=\mathbf{Y}_{\cdot n}/\sum_{d}Y_{dn}$(Srinivasan, Xue and Zhan, 2020). While we have used our notation $\mathbf{W}^{{\tiny\mbox{comp}}}$, it is important to note that CoDA is not formulated with any explicit notion of the system being measured. Said differently, CoDA does not distinguish between the observed data and the underlying system being studied. This is a critical difference between SRI and CoDA. We use the system $\mathbf{W}$ as a conceptual tool to define the true value of our estimand $\theta(\mathbf{W})$. This, combined with our probabilistic definition of compositional surveys, will allow us to investigate SRI and motivate our methods using classical inferential criteria (e.g., consistency and type-I error control). Lacking these structures, CoDA relies on three non-standard axioms to motivate its approach (Pawlowsky-Glahn, Egozcue and Tolosana-Delgado, 2015, Section 2.2). Chief among these is the principle of scale invariance, which states that any analysis of compositional data should be invariant to the arbitrary scale of the data. As a result of this axiom, CoDA restricts its attention to log-ratio estimands which are scale invariant: e.g., estimands of the form $\gamma=f(\eta_{1},\dots,\eta_{N})$ where $\eta_{n}=\log(\mathbf{W}_{in}/\mathbf{W}_{jn})$ (Lovell, Chua and McGrath, 2020). In contrast, SRI studies the estimation of scale reliant estimands.
Despite these critical differences, we choose terminology that is consistent with CoDA. For example, we use the term scale to refer to the sums $W^{{\tiny\mbox{tot}}}_{n}=\sum_{d}\mathbf{W}_{dn}$ of the system in the same way that CoDA uses the term to refer to the sums of the data $\sum_{d}\mathbf{Y}_{dn}$. We choose the term compositional survey to distinguish our probability model from the mathematical definition of compositional data while highlighting the intuitive relationship between them. Note our definition of compositional surveys does not require that the measurements $\mathbf{Y}_{\cdot n}$ are continuous or sum to a constant value. This is appealing as microbiome data lack information about the system scale yet are count vectors that do not sum to a constant value.
The implications of these differences between CoDA and SRI are discussed in later sections and summarized in Section 5.
2.4 The Continuum between SII and SRI
Identifiability is not a dichotomous state but rather a continuum based on the identification region of a parameter. This implies that there is a spectrum between SRI and SII. Here, we define this spectrum before showing how these definitions apply to inferential problems described in existing literature.
In the extreme case where ${\mbox{Id}}_{\theta}(P)$ is a singleton set, $\theta$ is identified, and the estimation problem is SII. At the opposite end where ${\mbox{Id}}_{\theta}(P)=\Theta$, $\theta$ is fully unidentified, and the estimation problem is Full SRI. Full SRI represents the extreme case where, due to the arbitrary scale of the data, all values of $\theta\in\Theta$ are consistent with the observed measurements. In between these two cases, where ${\mbox{Id}}_{\theta}(P)\subsetneq\Theta$, $\theta$ is partially identified, and the estimation problem is Partial SRI.
In Section 2.5, we show that the distinction between SII, Full SRI, and Partial SRI is directly related to the conditions under which certain types of inference (e.g., Frequentist versus Bayesian) are possible. Before providing those results, we first motivate that all three types of estimation problems are encountered within the literature. To do this, we describe the two relevant factors that determine whether a problem is SII, Full SRI, or Partial SRI. The factors are the degree to which the target estimand depends on scale and the size of the identification region of $W^{{\tiny\mbox{tot}}}$ at $P_{\mathbf{W}^{{\tiny\mbox{comp}}},W^{{\tiny\mbox{tot}}}}(\mathbf{Y})$.
2.4.1 Characterization of Target Estimands
Target estimands can be characterized based on the degree to which they depend on scale – whether they are scale invariant, partially scale reliant, or fully scale reliant. We define this characterization before providing examples of each type.
Let $\mathbf{W}^{{\tiny\mbox{comp}}}\times\mathcal{W}^{\tiny\mbox{tot}}$ denote the set of all systems that have a particular composition $\mathbf{W}^{{\tiny\mbox{comp}}}$. $\mathcal{W}^{\tiny\mbox{tot}}$ is the space of scale vectors, $\mathbb{Z}^{N+}$ if $W$ is integer valued or $\mathbb{R}^{N+}$ if continuous. Let $\theta(\mathbf{W}^{{\tiny\mbox{comp}}}\times\mathcal{W}^{\tiny\mbox{tot}}$) denote the image of this set with respect to the target estimand:
$\theta(\mathbf{W}^{{\tiny\mbox{comp}}}\times\mathcal{W}^{\tiny\mbox{tot}})=\{\theta\in\Theta:\exists\,\,W^{{\tiny\mbox{tot}}}\in\mathcal{W}^{\tiny\mbox{tot}},\theta(\mathbf{W}^{{\tiny\mbox{comp}}},W^{{\tiny\mbox{tot}}})=\theta\}.$
We define scale invariance/reliance based on the size of this image. If $\theta(\mathbf{W}^{{\tiny\mbox{comp}}}\times\mathcal{W}^{\tiny\mbox{tot}})$ is a singleton set, we say that $\theta$ is scale invariant at $\mathbf{W}^{{\tiny\mbox{comp}}}$; otherwise, it is scale reliant at $\mathbf{W}^{{\tiny\mbox{comp}}}$. We sub-categorize scale reliant estimands as fully scale reliant if $\mbox{span}(\theta(\mathbf{W}^{{\tiny\mbox{comp}}}\times\mathcal{W}^{\tiny\mbox{tot}}))=\Theta$ or partially scale reliant if $\mbox{span}(\theta(\mathbf{W}^{{\tiny\mbox{comp}}}\times\mathcal{W}^{\tiny\mbox{tot}}))\subsetneq\Theta$. Our definition reflects the following intuition. If $\theta$ is scale invariant, then its value is uniquely determined by $\mathbf{W}^{{\tiny\mbox{comp}}}$ alone and does not depend on $W^{{\tiny\mbox{tot}}}$. In contrast, if $\theta$ is partially scale reliant then $\mathbf{W}^{{\tiny\mbox{comp}}}$ limits the values that $\theta$ can take on but does not uniquely determine its value. If $\theta$ is fully scale reliant, then $\mathbf{W}^{{\tiny\mbox{comp}}}$ places no limits on the values $\theta$ can take on.
The scale reliance or scale invariance of a target estimand can depend on the composition $\mathbf{W}^{{\tiny\mbox{comp}}}$. For example, in a companion article, we have studied a set of non-parametric estimands (based on modified Kolmogorov-Smirnov statistics) which is scale invariant for some $\mathbf{W}^{{\tiny\mbox{comp}}}$ and scale reliant for others (McGovern, Nixon and Silverman, 2023). For simplicity, this manuscript only considers target estimands where their scale reliance or scale invariance holds for all $\mathbf{W}^{{\tiny\mbox{comp}}}$.
Example (Microbiome Data Analyses).
The LFC estimand described in Equation (2) is fully scale reliant. This can be established as follows. For any $\mathbf{W}^{{\tiny\mbox{comp}}}$ (and therefore any $\theta^{\tiny\mbox{comp}}_{d}\in(-\infty,\infty)$) and any $\theta_{d}\in(-\infty,\infty)$; there exists some $W^{{\tiny\mbox{tot}}}\in\mathcal{W}^{\tiny\mbox{tot}}$, and corresponding $\theta^{\tiny\mbox{tot}}\in(-\infty,\infty)$, such that $\theta_{d}=\theta^{\tiny\mbox{comp}}+\theta^{\tiny\mbox{tot}}$. That is, without restrictions on $\theta^{\tiny\mbox{tot}}\in(-\infty,\infty)$, fixing the value of $\theta^{\tiny\mbox{comp}}_{d}$ places no restrictions on the value of $\theta_{d}$.
At the opposite extreme, the types of log-ratio estimands purported by the field of compositional data analysis are scale invariant. For example, CoDA often analyzes log-ratio abundances between pairs of taxa $d_{i}$ and $d_{j}$ (Lovell et al., 2015):
$\eta_{n}=\log\frac{\mathbf{W}_{d_{i}n}}{\mathbf{W}_{d_{j}n}}.$ (4)
These quantities can be calculated from the system composition alone:
$\eta_{n}=\log\frac{\mathbf{W}_{d_{i}n}}{\mathbf{W}_{d_{j}n}}=\log\frac{\mathbf{W}^{{\tiny\mbox{comp}}}_{d_{i}n}}{\mathbf{W}^{{\tiny\mbox{comp}}}_{d_{j}n}}.$ (5)
In between these two extremes there are partially scale reliant estimands. For example, in the analyse of human gut microbiome data a common task is to identify pairs of taxa whose abundances covary on the log-scale:
$\Sigma_{d_{i},d_{j}}=\text{cov}(\log\mathbf{W}_{d_{i}\cdot},\log\mathbf{W}_{d_{j}\cdot}).$ (6)
These covariance estimands underlie the construction of ecological interaction networks (Kurtz et al., 2015; Friedman and Alm, 2012). In Supplementary Section B.1, we prove this estimand is partially scale reliant: $\mathbf{W}^{{\tiny\mbox{comp}}}$ provides a finite lower bound for this estimate but it is unbounded from above regardless of $\mathbf{W}^{{\tiny\mbox{comp}}}$.
2.4.2 Characterization of Compositional Surveys
We categorize compositional surveys as partial or full compositional surveys depending on whether $W^{{\tiny\mbox{tot}}}$ is fully or partially identified by the model $P_{\mathbf{W}^{{\tiny\mbox{comp}}},W^{{\tiny\mbox{tot}}}}(\mathbf{Y})$. That is, partial compositional surveys provide some information about the system scale ${\mbox{Id}}_{W^{{\tiny\mbox{tot}}}}(P)\subsetneq\mathcal{W}^{\tiny\mbox{tot}}$ whereas full compositional surveys provide no information ${\mbox{Id}}_{W^{{\tiny\mbox{tot}}}}(P)=\mathcal{W}^{\tiny\mbox{tot}}$. For brevity, this article focuses on full compositional surveys. Still, we highlight this distinction as we believe that further study of partial compositional surveys would be beneficial.
Example.
Authors who claim that microbiome data are compositional (e.g., Gloor et al. (2017)) are essentially stating that the data provides so little information about $W^{{\tiny\mbox{tot}}}$ that can be equated with simplicial vectors. We interpret such claims as a belief that the data comes from a full compositional survey. Empirical studies support this view illustrating that there is little relationship between the system scale and the scale of the data (Vandeputte et al., 2017; Props et al., 2017). While we expect that future researchers will identify specific cases where microbiome data does provide some degree of information about scale; for simplicity, this article models microbiome data as full compositional surveys.
Outside of microbiome studies there are clear cases where the data comes from partial compositional surveys. Consider political polling. Under standard conditions, a poll of 1000 voters in a town implies that the town must have at least 1000 voters. Still these data typically do not provide sufficient information to determine the true number of voters.
2.4.3 Theoretical Tools for Establishing SRI
In Supplemental Section B.2 we provide theoretical results which establish how estimands and compositional surveys combine to determine if an estimation problem is Full SRI, Partial SRI, or SII. We use those results to draw the following conclusions about estimation from full compositional surveys (e.g., microbiome data). LFC estimation (Eq 2, e.g., differential abundance analysis) is Full SRI. Covariance estimation (Eq. 6) is Partial SRI. Estimation of log-ratio quantities (e.g., Eq. 4) is SII.
2.5 Limits on Scale Reliant Inference
Having formally defined SRI as a problem of identifiability, we can prove fundamental limits on the types of inference that can be performed using data with arbitrary scale. We frame these limits in terms of common Frequentist and Bayesian inferential criteria. We show that there are strict limits on Frequentist inference but none on Bayesian inference. However, rigorous Bayesian inference requires an uncommon type of model called a Bayesian Partially Identifiable Model (PIM). We use the concept of unacknowledged bias to discuss the practical implication of these limits.
2.5.1 Limits on Frequentist Inference
We considers Frequentist inferential procedures motivated by canonical criteria such as consistency, unbiasedness, interval calibration, and type-I error control. The following four theorems follow from the definition of SRI as a problem of identifiability. Related impossibility proofs can be found in other fields that encounter partially identified models: such as econometrics (Dufour, 1997; Gabrielsen, 1978) or in the study of error-in-variable models (Gleser and Hwang, 1987). Proofs are provided in Supplementary Section B.3.
Theorem 2.2 (Consistency).
If ${\mbox{Id}}_{\theta}(P)$ is not a singleton set (Full or Partial SRI), then there does not exist a sequence of estimators for the target estimand $\theta$ that is consistent.
Theorem 2.3 (Unbiasedness).
If ${\mbox{Id}}_{\theta}(P)$ is not a singleton set (Full or Partial SRI), then there does not exist an estimator for the target estimand $\theta$ that is unbiased.
Theorem 2.4 (Calibration).
Suppose there exists $\theta\in\Theta$ and $\lambda\in\Lambda$ such that $Id_{\theta}(P)=\Theta$ (Full SRI). There does not exist a non-trivial [^1] See Supplementary Section B.3. confidence set $S(\mathbf{Y})$ calibrated at level $01-\alpha>$ 0.
Theorem 2.5 (Type-I Error Control).
Suppose there exists some $\theta\in\Theta$ and $\lambda\in\Lambda$ such that $Id_{\theta}(P)=\Theta$ (Full SRI). Then, for any $\alpha>0$, there does not exist a non-trivial [^2] See Supplementary Section B.3. level-$\alpha$ test of the null hypothesis $\theta\in\Theta_{0}\subsetneq\Theta$.
To the best of our knowledge, these are the first formal statements of what types of inferential criteria can or cannot be attained when analyzing data with arbitrary scale. These theorems provide formal justification for CoDA’s insistence on scale invariance, yet also show that existing arguments within the CoDA community are overly restrictive.
Consistent with CoDA’s arguments, these theorems show that the Frequentist ideal is not attainable for Full SRI problems. For example, Theorem 2, contradicts a growing body of literature which claims to have developed rigorous hypothesis tests for differential analysis of compositional data (Full SRI) that control Type-I error (Srinivasan, Xue and Zhan, 2020; Brill, Amir and Heller, 2022; Love, Huber and Anders, 2014; Fernandes et al., 2014). In Supplementary Section C, we review one paper that claims to have developed such a test and discuss how to reconcile those claims with our results.
Still, our results show that CoDA’s insistence on scale invariance goes to far. For example, our results show that non-trivial hypothesis tests could be developed for Partial SRI tasks. In fact, we have already done so in McGovern, Nixon and Silverman (2023). Moreover, if we look beyond Frequentist inference there are even more possibilities. In the next subsection, we show that Bayesian inference is possible for all forms of SRI.
2.6 Bayesian PIMs for SRI
The previous section established strict limits on Frequentist inference in the setting of SRI. That is, limits on inferential procedures motivated by common criteria such as consistency. In contrast, this section explains why there are no equivalent limits on Bayesian inference as standard Bayesian assumptions can be satisfied. There are two principle assumptions which we review below: the likelihood principle and the prior calibration principle. While we explain how these assumptions can be satisfied, we also show that doing so requires a specialized type of model called a Bayesian PIM.
Consider the problem of estimating LFCs for $D$ taxa ($\theta=(\theta_{1},\dots,\theta_{D})$) from a microbiome study $\mathbf{Y}$ (Full SRI). To satisfy the prior calibration principle, the prior must accurately represent prior beliefs about $\theta$. In most real-world scenarios we expect this would require a non-singular distribution $p(\theta)$, e.g., $\text{LogNormal}(\theta\mid\mu,\Sigma)$. Using Equation 3, such priors can reparameterized as $p(\theta^{\tiny\mbox{comp}},\theta^{\tiny\mbox{tot}})$ where $\theta^{\tiny\mbox{comp}}=(\theta^{\tiny\mbox{comp}}_{1},\dots,\theta^{\tiny\mbox{comp}}_{D})$ is a vector-valued function of $\mathbf{W}^{{\tiny\mbox{comp}}}$ and $\theta^{\tiny\mbox{tot}}$ is a scalar function of $W^{{\tiny\mbox{tot}}}$. To satisfy the likelihood principle, the chosen likelihood must accurately reflect the evidence provided by the observed data $\mathbf{Y}$ relevant to the model parameters $(\theta^{\tiny\mbox{comp}},\theta^{\tiny\mbox{tot}})$. However, since the observed data comes from a full compositional survey, satisfying the likelihood principle requires a likelihood invariant to $W^{{\tiny\mbox{tot}}}$:
$\mathcal{L}(\mathbf{W}^{{\tiny\mbox{comp}}},W^{{\tiny\mbox{tot}}};\mathbf{Y})\propto\mathcal{L}(\mathbf{W}^{{\tiny\mbox{comp}}};\mathbf{Y}).$
As an example, when analyzing multivariate count data, a product multinomial might be appropriate:
$\mathcal{L}(\mathbf{W}^{{\tiny\mbox{comp}}};\mathbf{Y})=\prod_{n}\text{Multinomial}(\mathbf{Y}_{\cdot n}\mid\mathbf{W}^{{\tiny\mbox{comp}}}_{\cdot n}).$
For LFC estimation, since $\theta^{\tiny\mbox{comp}}$ is solely a function of $\mathbf{W}^{{\tiny\mbox{comp}}}$ and $\theta^{\tiny\mbox{tot}}$ is solely a function of $W^{{\tiny\mbox{tot}}}$, this could be reparameterized as a likelihood of the form $\mathcal{L}(\theta^{\tiny\mbox{comp}},\lambda;\mathbf{Y})$ where $\lambda$ is some nuisance parameter that depends on $\mathbf{W}^{{\tiny\mbox{comp}}}$. For brevity, we omit $\lambda$ from future notation. This combination of non-singular prior $p(\theta^{\tiny\mbox{comp}},\theta^{\tiny\mbox{tot}})$ with a likelihood that is invariant to one of the prior parameters $\mathcal{L}(\theta^{\tiny\mbox{comp}};\mathbf{Y})$ is an example of a Bayesian PIM. A more general introduction to Bayesian PIMs can be found in Gustafson (2015).
Bayesian PIMs can have posteriors that demonstrate appealing, yet non-traditional behavior. Many traditional Bayesian models have posteriors which concentrate to a point-mass asymptotically (Schwartz, 1965). In SRI, this behavior is undesirable. From Theorem 2.3, any point estimator is biased and therefore any Bayesian posterior that asymptotically collapses to a point mass is problematic. Such methods are said to display an unacknowledged bias as they are biased yet fail to account for that bias in their estimation of uncertainty (Gustafson, 2015, pg. 35). Unlike those traditional methods, Bayesian PIMs often have non-zero asymptotic variance. More concretely, consider that the posterior for the above model factors as:
$p(\theta^{\tiny\mbox{comp}},\theta^{\tiny\mbox{tot}}\mid Y)$ $\propto\mathcal{L}(\theta^{\tiny\mbox{comp}};Y)p(\theta^{\tiny\mbox{comp}},\theta^{\tiny\mbox{tot}})$
$\propto\mathcal{L}(\theta^{\tiny\mbox{comp}};Y)p(\theta^{\tiny\mbox{comp}})p(\theta^{\tiny\mbox{tot}}\mid\theta^{\tiny\mbox{comp}})$
$\propto p(\theta^{\tiny\mbox{comp}}\mid Y)p(\theta^{\tiny\mbox{tot}}\mid\theta^{\tiny\mbox{comp}}).$ (7)
So long as the prior is non-singular, the posterior factor $p(\theta^{\tiny\mbox{tot}}\mid\theta^{\tiny\mbox{comp}})$ will not collapse, even asymptotically. The value of a posterior that maintains non-zero variance is straight-forward: if the posterior is biased, then only a posterior that has non-zero variance has any chance of covering the true value. It is for this reason that Bayesian PIMs are often said to display an acknowledged rather than unacknowledged bias. Reinforcing this terminology, one author has even shown that the posterior of Bayesian PIM’s can be interpreted as a probabilistic model for the bias of the posterior mean (Gustafson, 2015, pg. 35).
A major motivation of this work is to avoid unacknowledged bias. To underscore the danger, consider that a hallmark of unacknowledged bias is a Type-I or Type-II error rate that increases with more data (see Supplemental Figure 2 for visual explanation). This behavior is counter to standard statistical wisdom – where statistical procedures are supposed to improve with more data. For concreteness, in Supplemental Section D we discuss three common pitfalls encountered when developing Bayesian models for SRI including the use of Poisson or Negative Binomial likelihoods. That section discusses how each pitfall can lead to unacknowledged bias. In Section 4, we demonstrate the practical benefits of Bayesian PIMs for microbiome data analysis using both simulated and real data studies. Those studies demonstrate that common methods for microbiome differential abundance analysis demonstrate unacknowledged bias. They also show how Bayesian PIMs can drastically reduce Type-I and even Type-II error rates by acknowledging their own bias.
3 Scale Simulation Random Variables (SSRVs)
We now discuss practical solutions to SRI. We focus our attention on Bayesian PIM approaches as they can be applied to both Partial and Full SRI tasks. Two challenges confront the use of Bayesian PIMs for SRI: model specification and computational tractability.
Standard approaches to prior specification in Bayesian PIMs can lead to unreasonable assumptions about the system scale. For LFC estimation, a prior of the form $p(\theta)$ may make implicit assumptions about $\theta^{\tiny\mbox{tot}}$ though the conditional $p(\theta^{\tiny\mbox{tot}}\mid\theta^{\tiny\mbox{comp}})$. We demonstrate this pitfall in Section 4.2.
Beyond model specification, Bayesian PIMs can be challenging to infer. Returning to LFC estimation, note the posterior factor $p(\theta^{\tiny\mbox{tot}}\mid\theta^{\tiny\mbox{comp}})$ retains the variance of the prior while the variance of the factor $p(\theta^{\tiny\mbox{comp}}\mid\mathbf{Y})$ will typically decrease with increasing sample size. This leads to substantial anisotropy (i.e., $\mathrm{var}(\theta^{\tiny\mbox{comp}}\mid\mathbf{Y})<<\mathrm{var}(\theta^{\tiny\mbox{tot}}\mid\mathbf{Y})$) which poses computational challenges for traditional Bayesian inferential tools (Gustafson, 2015). Previous authors have suggested that tools like importance sampling could address these problems (Gustafson, 2015). In our experience, those methods work poorly for SRI where posteriors are often high-dimensional.
To address these challenges, we introduce Scale Simulation Random Variables (SSRVs), a modeling framework that can help specify and infer Bayesian PIMs for SRI.
3.1 Defining SSRVs
We start by defining SSRVs before discussing how they relate to Bayesian PIMs and address the aforementioned challenges. For simplicity, we limit our discussion to SSRVs for full compositional surveys. Generalizations to partial compositional surveys are straightforward yet may be more computationally challenging.
We define SSRVs as the combination of three components. The first is the target estimand, $\theta(\mathbf{W}^{{\tiny\mbox{comp}}},W^{{\tiny\mbox{tot}}})$. This estimand also defines the problem as SII versus SRI. The second is the measurement model, $p(\mathbf{W}^{{\tiny\mbox{comp}}}\mid\mathbf{Y})$. The measurement model represents posterior uncertainty in the system composition. The third component is a scale model, $p(W^{{\tiny\mbox{tot}}}\mid\mathbf{W}^{{\tiny\mbox{comp}}})$. The scale model is a prior distribution which represents uncertainty in the system scale and may be conditioned on the system composition. With these components, a SSRV is a random variable $\theta$ with probability distribution $p(\theta\mid\mathbf{Y})$ that can be defined via an integral transform.
Definition 3.1 (SSRVs: Integral Transform).
The following integral transformation defines a SSRV for $\theta$ as the transformation of the scale and measurement models through the target estimand:
[…]
The above integral transform suggests the following Monte-Carlo procedure to efficiently simulate SSRVs.
Data: $\mathbf{Y}$
Result: $S$ samples of $\theta$
for $s$ in $\{1,\dots,S\}$ do
       Sample $\mathbf{W}^{{\tiny\mbox{comp}}(s)}\sim p(\mathbf{Y})$
       Sample $W^{{\tiny\mbox{tot}}(s)}\sim p(\mathbf{W}^{{\tiny\mbox{comp}}(s)})$
       Calculate $\theta^{(s)}=\theta(\mathbf{W}^{{\tiny\mbox{comp}}(s)},W^{{\tiny\mbox{tot}}(s)})$
end for
Algorithm 1 Monte-Carlo Algorithm for Simulating SSRVs
Less formally, SSRVs are simulations of the target estimand that incorporate both uncertainty in the system composition and uncertainty in the system scale. This interpretation is echoed in the SSRV mean which is the expectation of the target estimand taken with respect to the measurement and scale models:
$\mathbb{E}[\theta\mid\mathbf{Y}]=\mathbb{E}_{p(W^{{\tiny\mbox{tot}}}\mid\mathbf{W}^{{\tiny\mbox{comp}}})}[\mathbb{E}_{p(\mathbf{W}^{{\tiny\mbox{comp}}}\mid\mathbf{Y})}[\theta(\mathbf{W}^{{\tiny\mbox{comp}}},W^{{\tiny\mbox{tot}}})]].$
3.2 Relationship to Bayesian PIMs
SSRVs are Bayesian PIMs which are specified using augmented variables $W^{{\tiny\mbox{tot}}}$ and $\mathbf{W}^{{\tiny\mbox{comp}}}$. Within the integral of Equation 8, the term $p(W^{{\tiny\mbox{tot}}}\mid\mathbf{W}^{{\tiny\mbox{comp}}})p(\mathbf{W}^{{\tiny\mbox{comp}}}\mid\mathbf{Y})$ is the posterior of a Bayesian PIM. Unlike the Bayesian PIMs discussed previously, the Bayesian PIM within the integral of Equation 8 uses $W^{{\tiny\mbox{tot}}}$ and $\mathbf{W}^{{\tiny\mbox{comp}}}$ as parameters rather than $\theta$. SSRVs then use the target estimand to transform this posterior over $\mathbf{W}^{{\tiny\mbox{comp}}}$ and $W^{{\tiny\mbox{tot}}}$ into a posterior over $\theta$. SSRVs are simply an alternative method of specifying Bayesian PIMs for SRI.
SSRVs mitigate the challenge of specifying and inferring Bayesian PIMs discussed above. Whereas priors placed directly on $p(\theta)$ may hide assumptions about scale within the conditional $p(\theta^{\tiny\mbox{tot}}\mid\theta^{\tiny\mbox{comp}})$, the scale model of an SSRV is that conditional. That is, assumptions about scale are an explicit part of SSRV specification. In Section 4.3 we highlight the benefit of this transparency using real and simulated data examples. SSRVs can also simplify posterior computations. Algorithm 1 only requires simulating from the measurement model and the scale model before combining those simulations in the target estimand. The measurement model is a standard (identified) Bayesian model, and the scale model is simply a prior. Sampling from both distributions is comparatively straightforward and avoids the anisotropy common with Bayesian PIMs.
4 Applications
This section considers real and simulated case studies of microbiome differential abundance analysis to illustrate the theory and methods developed in the prior sections. We start by reviewing the ALDEx2 (Fernandes et al., 2014) model which is a popular method for differential abundance analysis. We show that ALDEx2 can be thought of as an SSRV with an unrealistic scale model that underestimates prior uncertainty. Using real and simulated studies we show how this scale model leads to elevated rates of type-I and type-II errors in practice. Remarkably, we find similar patterns of false positives and false negatives with the distinct DESeq2 model (Love, Huber and Anders, 2014) suggesting our insights generalize beyond ALDEx2. We show how SSRVs address the pitfalls of these methods.
4.1 ALDEx2 is a SSRV with a Degenerate Scale Model
While there are many popular tools for differential abundance analysis, we highlight ALDEx2 (Fernandes et al., 2014) as it has a particularly simple mathematical form which aligns closely with SSRVs.
In Supplementary Section D.3, we present the ALDEx2 model as originally published in Fernandes et al. (2014). For brevity and clarity, we briefly review a simplified form of the ALDEx2 model which can be described in three steps. In the first step, posterior draws from independent multinomial-Dirichlet models are used to simulate uncertainty in the composition conditioned on the observed data. In the second step, those draws are normalized using the Centered Log-Ratio (CLR) transform which is a transform from the $D$-simplex ($\mathbb{S}^{D}$) to $\mathbb{R}^{D}$ defined as $\phi(x)=\log x-\mathbf{1}\operatorname*{mean}(\log x)$ for a vector $x\in\mathbb{S}^{D}$. In the third step, LFCs are calculated using the CLR-normalized relative abundances. The result is a distribution over LFCs. In fact, these three steps correspond to the measurement model, scale model, and target estimand of a SSRV.
The multinomial-Dirichlet samples imply a measurement model of the form:
$\mathbf{W}^{{\tiny\mbox{comp}}}_{\cdot n}\sim\text{Dirichlet}(\mathbf{Y}_{\cdot n}+\alpha)$
where $\alpha$ are the chosen Dirichlet prior parameters (default to $\alpha=0.5\cdot\mathbf{1}_{D}$ in ALDEx2). The target estimand is the LFC estimand (Eq. 2):
$\theta_{d}=\operatorname*{mean}_{n:x_{n}=1}\log\mathbf{W}_{dn}-\operatorname*{mean}_{n:x_{n}=0}\log\mathbf{W}_{dn}.$
Together, these two components suggest that the CLR normalization step represents some form of implicit scale model. This normalization can be written as:
$\log\mathbf{W}_{dn}=\log\mathbf{W}^{{\tiny\mbox{comp}}}_{dn}-\operatorname*{mean}\left(\log\mathbf{W}^{{\tiny\mbox{comp}}}_{\cdot n}\right).$
Comparing this to $\log\mathbf{W}_{dn}=\log\mathbf{W}^{{\tiny\mbox{comp}}}_{dn}+\log W^{{\tiny\mbox{tot}}}_{n}$ implied by Equation 3, it is clear that this normalization equates to a scale model of the form:
$\log W^{{\tiny\mbox{tot}}}_{n}\sim\delta[-\operatorname*{mean}\left(\log\mathbf{W}^{{\tiny\mbox{comp}}}_{\cdot n}\right)]$
where $\delta[x]$ denotes a point mass centered at the point $x$. This scale model has zero variance. That is, ALDEx2 does not account for any uncertainty in its implicit assumption regarding the unmeasured system scale. Errors in this assumption lead to biased estimates.
As an aside, we expect that similar unacknowledged bias due to errors in implicit assumptions regarding scale may be found in other fields. Normalizations are common when analyzing polling data (Topîrceanu and Precup, 2020) and in Supplementary Section F, we provide an example of an implicit assumption in a specific type of ecological survey.
4.2 ALDEx2’s Degenerate Scale Model Drives False Positives
We simulated a microbiome study comparing human gut microbiota in otherwise healthy individuals before and after treatment with a mild, narrow spectrum antibiotic whose true effect is to slightly decrease the abundance of 4 of the 20 simulated taxa (full simulation details provided in Supplementary Section G.1). Figure 1a shows the simulated true abundances ($\mathbf{W}$) for a set of $N=100$ study participants equally split between the pre- and post-antibiotic conditions. We simulated data $\mathbf{Y}$ with arbitrary scale by resampling the true abundances $\mathbf{W}$ with a multinomial distribution so the arbitrary scale of each sequenced sample $\mathbf{Y}_{\cdot n}$ has a total of $5,000$ total counts. We applied two popular DA models (DESeq2 (Love, Huber and Anders, 2014) and ALDEx2 (Fernandes et al., 2014)) to $\mathbf{Y}$ and, in Figure 1, show the pattern of true/false positives/negatives that result. In terms of Type-I Error control, both methods are worse than determining statistical significance by flipping a fair coin. Moreover, in Supplementary Figure 3 we show that the false discovery rate of both methods increases with increasing sample size – the hallmark of unacknowledged bias.
Figure 1: (a) Modeling results for DESeq2, ALDEx2, and a SSRV model. The Relaxed scale model is centered on the same geometric assumption as CLR normalization assumption but considers error in that assumption by using $\alpha=0.5$. b) Sensitivity analysis showing the estimated standardized log-fold change (LFC) for each taxon under the Relaxed model as a function of $\alpha$. The average and standard deviation of log fold change was computed across Monte Carlo samples, and a standardized value was computed for each entity by dividing the mean LFC by its corresponding standard deviation. A black line signifies that the estimated effect size was statistically different from zero at 95% confidence. The four taxa that were truly differentially expressed are labeled (Taxa 3, 4, 15, and 20).
As described in the prior section, we expect that the implicit scale model used by ALDEx2 underestimates a researcher’s true uncertainty in the system scale leading to an unacknowledged bias and driving the false positives seen in Figure 1. To evaluate this hypothesis, we created a relaxation of the ALDEx2 model replacing the degenerate (point-mass) scale model with the following relaxation: $\log W^{{\tiny\mbox{tot}}}_{n}\sim N(-\operatorname*{mean}\left(\log\mathbf{W}^{{\tiny\mbox{comp}}}_{\cdot n}\right),\alpha^{2})$.
Our relaxation is sufficient to remove all of ALDEx2’s false positives and retain most if not all of the true positives (Figure 1a). Additionally, our relaxation addresses the unacknowledged bias and does not lead to the same asymptotic increase in false discovery rates as seen with ALDEx2 and DESeq2 (Supplemental Figure 3). Moreover, sensitivity analysis shows that our results are largely insensitive to the choice of $\alpha$: even small amounts of uncertainty are sufficient to remove the majority of false positives at the cost of just a single false negative (taxon 3; Figure 1b). Only when $\alpha$ is chosen to be implausibly large ($\alpha>1.4$) does the relaxed ALDEx2 model fail to find true positives. Considering that the CLR estimate of scale corresponds to $\hat{\theta}^{\perp}_{CLR}=0.03$ a value of $\alpha>1.4$ reflects a greater than 5% probability that the log-fold-change of scales, $\theta^{\perp}$, is outside of the range $[\exp^{-1.96\times 1.4^{2}+\hat{\theta}^{\perp}_{CLR}},\exp^{1.96\times 1.4^{2}+\hat{\theta}^{\perp}_{CLR}}]=[0.02,48.02]$; which is implausible for a narrow spectrum antibiotic. In summary, our results indicate that the CLR normalization in ALDEx2 drives its observed false positives and that these can be mitigated by considering potential errors in assumptions about scale. The observed pattern of false positives within DESeq2 suggests similar effects may be at play within that model even though it uses a different median-of-ratios normalization.
4.3 Application to a novel 16S rRNA-Seq dataset
Although artificial gut systems (bioreactors) are seeded with human feces and often used as a model for host-associated microbiota, many bioreactors undergo a two week equilabration period in which the inoculated communities diverge from those seen in human hosts and adapt to their new ex vivo environment. To evaluate the limits of these model systems, we ran and collected data from 8 bioreactors over the course of the first two weeks after inoculation of which 6 were used for analyses due to experimental constraints (see Supplemental Section G.2). Each bioreactor was seeded with 100mL of fecal slurry and maintained at a constant 400mL of growth media over the course of the experiment. Our primary interest was in determining which of the microbial taxa present in the innoculum were able to grow in the bioreactors. We quantify this interest using the differential abundance target estimator given in Equation (2) with $x_{n}=0$ denoting day 1 of the experiment and $x_{n}=1$ denoting day 14. In this case, we were primarily interested in the de facto total scale, i.e., we wanted to identify taxa that increased or decreased in total cellular number over the two week period.
This data is ideal for our study as our experimental system allowed us to obtain high-quality measurements of scale in this bioreactors. Because each vessel was homogeneously mixed, microbial concentration measured from a sample of the vessel via flow cytometry was presumably identical to microbial concentration within the vessel. Moreover, because each vessel was maintained at a constant 400mL volume throughout the course of the experiment, changes in microbial concentration between vessels or between time-points were proportional to changes in total microbial load. Flow cytometry measurements were obtained to measure total cellular concentration. These measurements were taken in triplicate for each 16S rRNA sequencing measurement to allow us to quantify measurement noise. Further experimental details are presented in Supplemental Section G.2.
Five models were compared on the resultant data: one specified as a Bayesian PIM, two specified as SSRVs, ALDEx2, and DESeq2. We used the following multinomial log-normal Bayesian PIM, which is an extension of the Bayesian multinomial logistic-normal models that have become popular for analysis of microbiota survey data (Silverman et al., 2022):
$\mathbf{Y}_{\cdot n}$ $\sim\text{Multinomial}(\mathbf{W}^{{\tiny\mbox{comp}}}_{\cdot n})$ (9)
$\mathbf{W}_{\cdot n}$ $=\exp({\psi_{\cdot n}})$ (10)
$\psi_{\cdot n}$ $\sim N(\mathbf{B}X_{\cdot n},\boldsymbol{\Omega})$ (11)
with priors $\mathbf{B}\sim N(\mathbf{M},\boldsymbol{\Omega},\boldsymbol{\Gamma})$ and $\boldsymbol{\Omega}\sim IW(u,\boldsymbol{\Xi})$. This is a Bayesian PIM since the multinomial likelihood is only parameterized in terms of $\mathbf{W}^{{\tiny\mbox{comp}}}$ whereas prior is defined over both $\mathbf{W}^{{\tiny\mbox{comp}}}$ and $W^{{\tiny\mbox{tot}}}$. We chose the prior hyper-parameters to echo standard defaults we have observed in the literature, i.e., we chose $\mathbf{M}=0$ and choose $\mathbf{\Gamma}$ and $\mathbf{\Xi}$ to be diagonal matrices (see Supplementary Section G.2 for full prior specifications).
There are two critical limitations of this Bayesian PIM. First, it is computationally infeasible to infer this model directly. Instead, to obtain posterior samples of the parameter $\mathbf{B}$, we had to represent this model as an SSRV and sample $\mathbf{B}$ using Algorithm 1. Supplementary Section H derives the SSRV representation of this Bayesian PIM and shows that the measurement model is a Bayesian multinomial logistic-normal model and the scale model is a log-normal linear model. The second, more important, limitation of this model is that the assumptions about scale are not transparent. In fact, this seemingly innocuous and standard prior specification actually implies strong, biologically implausible assumptions: implicit in this prior is a non-trivial conditional relationship of the form $p(\theta^{\tiny\mbox{tot}}\mid\theta^{\tiny\mbox{comp}})$. This leads to a posterior for $\theta^{\tiny\mbox{tot}}$ which reflects a strong, biologically unreasonable belief that total microbial load is higher on day 1 than on day 14 (see Supplementary Section G.2). Below we show this leads to unacknowledged bias which presents as elevated Type-I and Type-II error rates.
To illustrate how SSRVs can enhance model specification and ease prior specification, we designed two SSRVs using the same measurement model as the above Bayesian PIM but with two different scale models. Letting $n\in\{1,\dots,N\}$ denote a given vessel at a given time-point, these two scale models were both log-normal models of the form :
$\log W^{{\tiny\mbox{tot}}}_{n}\sim N(c+\mu_{n},\alpha_{n}^{2})$ (12)
where $c$ is a constant and $\mu$ is subject to the constraint $\sum_{n=1}^{N}\mu_{n}=0$. Note that the LFC target estimand given by Equation 2 is invariant to the value of $c$ and so we can set $c=0$ with no loss of generality.
We created the Design scale model based only on prior knowledge of the experimental design. Since we know that each vessel contained 400mL of growth media and the innoculum consisted of 100 mL of fecal slurry, we expect that the total microbial load would, on average, increase from the day 1 to day 14 by a factor of approximately 4. We encapsulated these beliefs in the scale model
$\frac{f(W^{{\tiny\mbox{tot}}}_{{j},\text{t=1}})}{f(W^{{\tiny\mbox{tot}}}_{{j},\text{t=14}})}\sim\mathrm{LogNormal}(\bar{d},\tau^{2})$
where $\bar{d}=\log(100/400)$ and $\tau=0.5$.
Finally, we created the Flow scale model built from the flow-cytometry measurements. In this model $\mu_{n}$ and $\alpha^{2}_{n}$ were estimated as mean and variance our flow cytometry measurements on the log scale. We consider the SSRV created using this scale model our gold standard against which we define false positives. For concreteness, we define a result as a false positive in two cases: first, that the result is significant in a given model but not in the gold standard and second, that the result is significant in both the model and the gold standard, but the two models disagree on sign.
We applied the Bayesian PIM, the two SSRVs, ALDEx2, and DESeq2 to the artificial gut data; the results are summarized in Figure 2. As before we find that ALDEx2 and DESeq2 have high inferential error rates. However, unlike the prior study where these were primarily Type-I errors, here we observe far more Type-II errors owing to the fact that most taxa are actually increasing in abundance as the community grows to occupy the larger vessels and this study has a relatively small sample size ($N=12$). Still we find the hallmark of unacknowledged bias: using resampling, Supplemental Figure 5 shows that ALDEx2 and DESeq2 display increasing rates of false discoveries with increasing sample size while consistently displaying Type-II error rates far in excess of the Design model. For example with just 20 vessels, the FDR of ALDEx2 and DESeq2 are approximately 30% and 15% respectively. In contrast, even at 50 vessels, the Design model controls FDR at a nominal 5%. As expected based on our investigation of the Bayesian PIM prior, the Bayesian PIM model displays a false discovery rate of 45% which further increases with more data (Supplemental Figure 5).
Finally, Figure 3 shows that the major conclusions of this analysis, namely the remarkably performance improvement of the Design model compared to alternatives is largely insensitive to the choice of the parameter $\tau$. So long as $\tau$ is greater than 0.35, the FDR of the Design model will be less than 3.5%. Moreover, so long as we choose $\tau$ less than 5, the Type-II error rate of the design-based SSRV will be less than that of either DESeq2 or ALDEx2. A value of $\tau\geq 5$ would require we accept, with 2.5% probability, that the microbial load at day 14 was at least $4\exp(1.96*5^{2})\approx 7.6\times 10^{21}$ times greater than at day 1. Even if there was only one microbe at Day 1, this is not physically possible: as bacterium weights approximately $10^{-12}$ grams, this would require the vessels at day 14 to weight more than $10^{6}$ kilograms.
Figure 2: Comparison of results from ALDEx2, DESeq2, the Bayesian PIM, and two SSRVs using real data from an ex vivo artificial gut study. “GS” denotes the gold-standard model against which true/false positives/negatives are defined. Our gold standard (the SSRV using the Flow scale model) is the only model that could use the replicate flow-cytometry measurements collected in this study. For brevity, taxa that were identified as differentially abundant with respect to any of the 5 models are not shown (based on p-value less than 0.05 or posterior 95% credible interval not covering zero).
Figure 3: Sensitivity analysis of the Design SSRV as a function of the parameter $\tau$ in the Design scale model. The mean, standard deviation, and 95% credible interval were computed using the posterior distribution of log-fold change for each entity. The standardized log-fold change (LFC) was calculated by dividing the mean log-fold change by the corresponding estimate for standard deviation. The standardized LFC as a function of $\tau$ is shown for each taxon in the artificial gut study and is colored black if the corresponding credible interval excluded zero and grey otherwise. As an illustration, we highlight the 5 taxa that are the most clearly differentially abundant in the sense that their credible sets exclude zero over the largest range of $\tau$. For comparison, we depict the value of $\tau$ used in Figure 2 ($\tau=0.5$) as a black dotted line.
5 Discussion
In this article, we have framed the challenge of data that lacks scale information as an estimation problem involving partially identified models. Through this formulation, we have been able to provide the first formal limits on inference with these data. Our results are consistent with the CoDA perspective: they show that care is needed as spurious scientific conclusions can be reached due to the lack of information about the system scale contained in these data. However, unlike the CoDA perspective, our results argue that a dogmatic restriction to scale invariance is overly prohibitive. Consider that, even if data is compositional (i.e., a full compositional survey with simplex valued measurements), Bayesian PIMs can be used to perform SRI while satisfying Bayesian ideals. In fact, hypothesis tests and confidence intervals which satisfy Frequentist ideals can be constructed from compositional data if the estimand is partially scale reliant. Overall, we argue that SRI is possible yet requires new types of statistical methods that can account for the limited information contained in the observed data.
Beyond discussing limits on inference, this article introduces the class of Bayesian PIMs for SRI. There is an interesting connection between the SSRV formulation of Bayesian PIMs and adversarial robustness. Adversarial methods are well known in the machine learning literature as a way of enhancing the robustness or stability (Yu, 2020) of predictions. For example, within computer vision, noise is sometimes added to images in order to ensure that trained models are stable and are not sensitive to slight variation when applied out-of-sample (Santos and Papa, 2022). Due to concerns about rigor and reproducibility in science, some authors have advocated that adversarial perturbations can be applied to ensure robustness of statistical inference (Yu, 2020). Within SSRVs, the use of the scale model can be viewed as a type of adversarial perturbation: for each sample from the SSRV we apply a different random perturbation sampled from the scale model. One way of understanding the utility of Bayesian PIMs (and SSRVs) is that they involve a form of adversarial robustness to perturbations of scale.
In this article, we have focused on Bayesian approaches to SRI as these methods, unlike Frequentist methods, remain valid even in the setting of Full SRI. To clarify the problem with Frequentist methods, consider the problem of confidence estimation. Intuitively, a level-$\alpha$ confidence set for an estimand $\theta$ should, with probability $1-\alpha$, cover the identification region ${\mbox{Id}}_{\theta}(P)$. The problem is that, for Full SRI, ${\mbox{Id}}_{\theta}(P)=\Theta$ and therefore any level-$\alpha$ confidence set must be uselessly large. Yet, for Partial SRI, useful confidence estimation (and hypothesis testing) may be possible as ${\mbox{Id}}_{\theta}(P)\subsetneq\Theta$. For example, we recently proved that useful and robust hypothesis testing is possible for a particular partial SRI problem called differential set analysis (McGovern, Nixon and Silverman, 2023). Overall, we believe that Frequentist SRI for Partial SRI problems is an intriguing and likely fruitful area for future research. In fact, there is already a rich literature on Frequentist confidence estimation for partially identified models can likely be adapted to the task. We refer interested readers to Imbens and Manski (2004) and Stoye (2009) for a thorough review of those methods.
For simplicity, we have focused on compositional surveys which are a subset of data with arbitrary scale where the system composition is identifiable. Yet in many practical applications, we expect that the system composition may not be identifiable as the measurement process may involve unobserved compositional biases. For example, in the 16S rRNA microbiome surveys, the PCR amplification process can lead to an over-representation of some taxa compared to others in the sequenced pool (Silverman et al., 2021). If compositional bias is substantial, then it may tighten the limits on inference further restricting the types of inference that can be performed rigorously with these data. Just as Bayesian PIMs can be used to address arbitrary scale, we expect these same models may be adapted to provide robustness to compositional bias. Overall, we believe that the study of compositional bias represents a fruitful area for future research.
{acks}
[Data Access and Availability] All data and R scripts necessary to replicate all analyses in the manuscript and supplement are located at https://github.com/michellepistner/ScaleSimulation.
{acks}
[Acknowledgments] We thank Rachel Silverman, Juan Jose Egozcue, Vera Pawlosky-Glahn, Gregory Gloor, Michael Love, and James Morton for their manuscript comments.
{funding}
JDS and MPN were supported in part by NIH 1R01GM148972-01. LAD acknowledges support from the the Global Probiotics Council, a Searle Scholars Award, the Hartwell Foundation, an Alfred P. Sloan Research Fellowship, the Translational Research Institute through Cooperative Agreement NNX16AO69A, the Damon Runyon Cancer Research Foundation, and NIH 1R01DK116187-01. SM would like to acknowledge the support of grants NSF IIS-1546331, NSF DMS-1418261, NSF IIS-1320357, NSF DMS-1045153, and NSF DMS-1613261.
[…]